### v1

Abstract
We study decision making under risk when perception may be menu-dependent. Behavior is modeled as the outcome of a small library of psychologically interpretable, menu-specific rules that transform each objective menu into a perceived one. At each menu, the applied rule must make the realized choice a strict improvement under a dominance benchmark on perceived lotteries. We introduce the Maximum Rule Concentration Index (MRCI), the maximal Herfindahl-Hirschman concentration of rule shares over all locally admissible assignments, and diagnostics that distinguish rules that unify behavior across many menus from rules that mainly act as substitutes. We provide a MIQP formulation, a scalable heuristic, and a finite-sample permutation test of excess concentration relative to a menu-independent random-choice benchmark. Applied to the CPC18 dataset ($N=686$ subjects, each making 500500-$700$ repeated binary lottery choices), the mean MRCI is $0.545$, and $64.1\%$ of subjects reject random choice at the $1\%$ level. Concentration gains are primarily driven by modal-payoff focusing, salience-thinking, and regret-based comparisons.
JEL D91, D11, D81
Keywords: Decision Theory, Revealed Preference, Behavioral Economics.
1 Introduction
Choices under risk are famously patterned. In laboratories and in the field, people violate monotonicity, reverse rankings when lotteries are mixed with a common consequence, and display menu-dependent behavior that resists any single, stable ordering over objective lotteries. Over the past decades, the literature has responded with an impressive array of models - probability weighting, reference dependence, salience, regret, disappointment, limited attention - each offering a psychologically plausible channel through which the same objective menu may be perceived and evaluated differently. [^1] See, among many others, kahneman_tversky1979; tversky_kahneman1992; prelec1998 on probability weighting; KoszegiRabin2006 on reference dependence; bordalo2012_risk on salience in risky choice; LoomesSugden1986; bell85; Gul1991 on regret and disappointment; and sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014 on limited attention and consideration. This richness has greatly expanded our understanding of risky choice, but it also raises a complementary empirical question. When perception may vary with the menu, many distinct behavioral mechanisms can potentially explain the same set of risky decisions. This motivates a different emphasis: in a large dataset, which behavioral mechanisms are actually carrying the explanatory burden, how many distinct mechanisms are needed to organize behavior, and how stable those conclusions are as the set of candidate explanations is enlarged.
In this paper, we study binary risky choice through the lens of menu-specific decision rules. We posit a finite library of psychologically interpretable transformations - worst-case focusing (max-min), coarse probability distortion, salience-based focusing on the most contrasting state, regret- and disappointment-based comparisons, and one-sided attention that map each objective binary menu into a perceived one. The analyst does not observe these perceptions; she observes only the realized choices. Our identifying restriction is therefore local: at each menu, the rule assigned to that menu must render the chosen perceived lottery a strict improvement under a First Order Stochastic Dominance. This menu-level discipline has a sharp consequence: any locally admissible rule profile induces a perceived dataset that is automatically cycle-free in revealed-preference terms. Thus, the empirical content of the framework is not obtained by imposing global transitivity on observed choices. It comes from asking how parsimoniously one can assign rules across menus while preserving local dominance-based justification.
We operationalize parsimony with a new statistic, the Maximum Rule Concentration Index (MRCI). For a given dataset and rule library, we consider all menu-by-menu rule assignments that satisfy the local discipline. Each admissible assignment induces an empirical distribution of rule usage across menus. The MRCI is the largest achievable concentration of that distribution, measured by the Herfindahl-Hirschman index of rule shares: it is high when a dataset can be explained by repeatedly invoking the same rule across many menus, and low when any admissible explanation must spread weight across multiple rules. We complement the index with two rule-level diagnostics designed to answer the first paragraph’s question of “which mechanisms are doing the work.” The stability score records how often a rule remains necessary to attain maximal concentration when rules are deleted in different orders, while the concentration gain measures how much maximal concentration falls when a rule is removed. Together, these objects distinguish rules that genuinely unify behavior across menus from rules that only matter as interchangeable substitutes.
Computing the MRCI is a combinatorial optimization problem: we must assign to each menu a locally admissible rule so as to maximize a convex concentration functional. We provide an exact mixed-integer quadratic programming formulation in the Appendix, but it scales exponentially with the number of observations. We therefore develop a scalable heuristic that exploits a simple implication of convexity: concentration-maximizing solutions tend to load heavily on a small subset of high-coverage rules, so the search can focus on rule orderings that favor such high-coverage candidates. Our algorithm combines (i) a preprocessing step that determines local admissibility rule-by-rule and menu-by-menu, (ii) a greedy assignment pass for a given rule ordering, and (iii) random restarts over many orderings. A benchmarking exercise against the MIQP on subsampled problems shows that the heuristic is accurate, with a mean absolute error below $0.001$, more than $96\%$ of estimates within a $1\%$ tolerance, and it runs in negligible time.
We also develop inference. High concentration might be mechanically attainable by chance, even in environments where the admissibility constraints are nontrivial. To address this, we propose a finite-sample permutation test in the spirit of cherchye2025. The test is grounded in a Random Rule Model interpretation that is specific to binary menus. Because each choice problem has two labeled alternatives (left/right) and the library includes the two one-sided attention defaults (“attend to the left” and “attend to the right”), any realized sequence of left/right choices can be replicated by random switching between these two rules. [^2] This is similar to treatments of the power of nonparametric tests of preference maximization.See bronars and beatty2011. Under the null that choices are exchangeable across menus (menu-independent randomness), the conditional distribution of the MRCI given the realized number of left versus right choices is obtained by permuting the left/right indicators across fixed menus. We compare the observed MRCI to the conditional randomization distribution. Rejection means that the observed maximal concentration is too large to be plausibly generated by menu-independent random switching, providing statistical content to the notion that “few-rule behavior” is genuinely present in the data.
We apply the approach to the CPC18 dataset (plonsky2017_cpc18), which provides rich within-subject repetition over a common set of binary lottery problems designed to elicit canonical anomalies. Since our dominance benchmark and rule library are defined over lotteries with known probabilities, we restrict attention to the risky (non-ambiguous) problems. The resulting sample comprises $N=686$ subjects; each subject faces between $20$ and $28$ distinct binary lottery problems, each repeated for $25$ trials, yielding 500500-$700$ decisions per subject. The results show that behavior is typically far from a one-rule description but also far from requiring a large portfolio of rules. The average MRCI is $0.545$ (s.d. $0.050$), corresponding to an effective number of rules $N_{\mathrm{eff}}=1/\operatorname{MRCI}\simeq 1.85$.
Our inference procedure shows that this concentration is not merely mechanical. Implementing the permutation test, we find that $64.1\%$ of subjects reject the random-choice benchmark at the $1\%$ level. In other words, for a large share of subjects, the observed degree of concentration is too high to be plausibly generated by a menu-independent random process once we condition on the realized frequency of left/right choices.
The rule-importance diagnostics then identify which mechanisms account for this concentration. We find that MAP (modal-payoff focusing) is the main unifier on average, accounting for about $17\%$ of attainable concentration, with SAL (salience-based focusing) next at roughly $7\%$; regret-based comparisons (REG) and the extremum heuristics (MMx/MMn) contribute smaller but meaningful shares (on the order of $4$-$5\%$ each). Stability is more evenly distributed across the active rules: while MAP is most often indispensable (with a stability score $40\%$), REG and SAL also remain necessary in a nontrivial fraction of cases (roughly $25\%$ and $20\%$), indicating substantial substitution across mechanisms even when MAP drives the average gain. By contrast, PW and Id have negligible importance. This might be because these rules often fail to induce strict FSD improvements menu-by-menu. DIS appears genuinely non-operative in CPC18, in the sense that it almost never affects maximal concentration and is rarely needed to sustain it.
Our framework speaks to three strands of the literature. First, it complements the large decision-theoretic literature that explains patterned risky choice by enriching preferences or perception, including probability weighting, reference dependence, salience, regret, disappointment, and limited attention (kahneman_tversky1979; tversky_kahneman1992; prelec1998; KoszegiRabin2006; bordalo2012_risk; LoomesSugden1986; bell85; Gul1991; masatlioglu2012; manzini_mariotti2014; caplin_dean2015). We complement this literature, as we provide a nonparametric way to summarize how much structure is needed to render behavior locally defensible, and which broad mechanisms are most consistent with the data once they compete on a common footing. This perspective differs from approaches that prioritize flexible prediction with large, high-dimensional models or machine-learning toolkits, including recent work that achieves high predictive accuracy in risky-choice prediction settings, including erev2011, erev_ert_plonsky_2017, plonsky2016, or peterson2021.
Second, the paper connects to the growing literature that models choice through procedures and simple rules rather than through a single stable preference ordering. This tradition includes bounded rationality and heuristic choice (simon1955; tversky_kahneman1974; Kahneman1982; rubinstein88), “rule rationality” arguments emphasizing the efficiency of simple rules in natural environments and their fragility in contrived ones (aumann2008; aumann2019), simplicity in choices under risk (enke2023; fudenberg2022; puri2025), automata-based complexity constraints (salant2011menu), and procedural models of consideration, shortlisting, and categorization (salant2006; salant_rubinstein2008; manzini2007; manzini2012; cherepano2013), as well as recent experimental work that studies procedure choice directly (halevy2024). We are deliberately agnostic about why a given decision maker uses a particular heuristic or when a rule is optimal; our focus is on identification in standard risky-choice datasets, where rules are latent and must be inferred from choices.
Third, the paper contributes to the revealed preference literature. Classic revealed-preference theory, from samuelson1938note onward, makes cyclical consistency central by treating global rationalizability as the core testable implication of utility maximization (afriat1967; varian1982; houthakker1950). [^3] Using an experimental design that extends the canonical budget-line paradigm of syngjoo2007 to a three-dimensional risky-choice environment, dembo2021 find that departures from expected-utility theory go beyond violations of independence, and that many prominent non-EUT alternatives - which nonetheless preserve first-order stochastic dominance - do not fully account for these patterns. polisson2020 develops a nonparametric revealed-preference approach for testing a broad class of choice under risk models using data from budget-line experiments. In our framework, discipline comes instead from a local dominance-based justification applied after rule-induced transformations. Under first-order stochastic dominance, this local discipline implies a strong form of cyclical consistency in the induced perceived dataset. Methodologically, we also complement the revealed-preference computation and inference literature that relies on mixed-integer programming to compute indices and test statistics (HEUFER201587; Demuynck2023) by providing a scalable heuristic for our concentration object, together with a permutation test similar to cherchye2025.
The rest of the paper proceeds as follows. Section 2 introduces the choice environment, the rule library, and formalizes admissible rule profiles. Section 3 develop several examples, and Section 4 defines the MRCI and the rule-importance diagnostics. Section 5 develops a scalable computational procedure. Section 6 provides the permutation test. Section 7 reports the empirical results, and Section 8 concludes.
2 The Model
We consider a decision maker facing a sequence of binary choices under risk. Let $X$ denote the set of all finite-support lotteries over real monetary outcomes. A lottery $L\in X$ is represented by a pair $(z,\pi)$, where $z=(z_{1}<\dots<z_{m})\in\mathbb{R}^{m}$ are the ordered distinct prizes and $\pi\in\Delta^{m-1}$ is the associated probability distribution.
[…]
$L^{1}\triangleright L^{2}\quad\Longleftrightarrow\quad L^{1}\trianglerighteq L^{2}\ \text{ and not }(L^{2}\trianglerighteq L^{1}).$
Throughout this paper, we adopt First-Order Stochastic Dominance (FSD) as the dominance criterion.
Definition 1 (First-Order Stochastic Dominance).
Let $L^{1},L^{2}\in X$. Let $\mathrm{supp}(L^{1})\cup\mathrm{supp}(L^{2})=\{z_{1}<\dots<z_{M}\}$ be the ordered union of their supports. For $\ell\in\{1,2\}$, let $\pi^{\ell}_{k}\ :=\ \Pr_{L^{\ell}}(z_{k})$ with the convention $\pi^{\ell}_{k}=0$ when $z_{k}\notin\mathrm{supp}(L^{\ell})$, and define the right-tail cumulative sums $S^{\ell}_{i}\ :=\ \sum_{k=i}^{M}\pi^{\ell}_{k}$, $i=1,\dots,m$. We say that $L^{1}\geq_{\mathrm{FSD}}L^{2}$ if and only if $S^{1}_{i}\geq S^{2}_{i}$ for all $i=1,\dots,M$. We denote the strict part by $_{\mathrm{FSD}}$.
Acts and States. For some of our decision rules introduced later, it is convenient to think of each lottery as induced by a state-contingent act. Formally, a choice problem can equivalently be described by a finite state space $S=\{1,\dots,n\}$, objective state probabilities $(p_{s})_{s\in S}$, and, for each menu $A_{t}=\{L_{t}^{1},L_{t}^{2}\}$, two acts corresponding to the two lotteries:
$L_{t}^{i}\equiv\big(u_{t}^{i}(s)\big)_{s\in S},\qquad u_{t}^{i}(s)\in\mathbb{R}.$
Given $(p_{s})_{s\in S}$, the act $L_{t}^{i}$ induces a lottery representation $(z_{t}^{i},\pi_{t}^{i})\in X$ by grouping equal prizes and assigning to each distinct payoff its total probability:
$\pi^{i}_{t,k}\;=\;\sum_{s\in S:\,u_{t}^{i}(s)=z^{i}_{t,k}}p_{s}.$
Hence our lottery representation is without loss of generality for the properties we study. We will use the state-space view only when defining several rules, and otherwise work with the simpler lottery notation introduced above.
2.1 Bounded Rationality
In this paper, we model behavior as the outcome of a small library of menu-dependent decision rules. Each decision rule transforms the objective menu into a perceived one, after which the decision maker behaves as if she were choosing according to a dominance-respecting ordering on perceived lotteries.
Formally, let $\mathcal{F}$ be a finite collection of decision rules. Each rule
[…]
The rule library is deliberately simple and nonparametric. It collects ten parameter-free rules that span four broad mechanisms emphasized in the bounded-rationality and behavioral-decision literatures: outcome simplification, probability distortion, context-dependent comparisons, and limited attention or consideration. [^4] On simple heuristics and “fast and frugal” rules see, e.g., tversky_kahneman1974; gigerenzer_gaissmaier2011; brandstatter_et_al2006. On probability weighting and cumulative prospect theory, see kahneman_tversky1979; tversky_kahneman1992; prelec1998. On salience, regret and disappointment, see bordalo2012_risk, LoomesSugden1986; bell85; Gul1991, and KoszegiRabin2006. On limited attention and consideration, see, among others, sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014; caplin_dean2015. We now define each rule formally.
Identity (Id).
The decision maker perceives each lottery as it is:
$\pi_{\textsc{Id}}(L^{i};L^{j})=L^{i}.$
[…]
Outcome Simplification Rules
MaxMin (MMn).
The decision maker replaces $L^{i}$ by its sure minimum payoff:
$\pi_{\textsc{MMn}}(L^{i};L^{j})=\big(z^{i}_{1}\ \text{with prob.\ }1\big).$
This “safety-first” rule focuses exclusively on worst-case outcomes.
MinMax (MMx).
The decision maker replaces $L^{i}$ by its sure maximum payoff:
$\pi_{\textsc{MMx}}(L^{i};L^{j})=\big(z^{i}_{m_{i}}\ \text{with prob.\ }1\big).$
This rule is the bold counterpart of MaxMin: attention is placed entirely on best-case outcomes.
Max at $p$ (MAP).
The decision maker replaces $L^{i}$ by the sure modal payoff of $L^{i}$ (the payoff with highest probability). If several payoffs are tied for modal probability, we select the highest among the tied payoffs:
$k^{\star}(L^{i})\in\arg\max_{k\in\{1,\dots,m_{i}\}}\ \pi^{i}_{k},\qquad k^{\star}(L^{i})\in\arg\max\{\,z^{i}_{k}:\ \pi^{i}_{k}=\max_{q}\pi^{i}_{q}\,\},$
and define
$\pi_{\textsc{MAP}}(L^{i};L^{j})=\big(z^{i}_{k^{\star}(L^{i})}\ \text{with prob.\ }1\big).$
Here, only the modal consequence is retained and tails are neglected. (This is the “Max at pp” heuristic: pick the payoff at the most likely state.)
MaxMin, MinMax, and Max at $p$ all apply the same “score” functional to each lottery (minimum payoff, maximum payoff, or modal payoff) and compare lotteries on that scalar feature.
Probability Distortion Rule
Probability Weighting (PW).
The PW rule preserves outcomes and distorts probabilities via a coarse, S-shaped map. For $p\in[0,1]$, define
$\tilde{p}(p)\ :=\begin{cases}0.2,&\text{if }p\leq 0.2,\\[3.00003pt] p,&\text{if }0.2<p<0.8,\\[3.00003pt] 0.8,&\text{if }p\geq 0.8.\end{cases}$
Given a lottery $L^{i}=(z^{i},\pi^{i})$, we distort each objective probability $\pi^{i}_{k}$ to $\tilde{p}(\pi^{i}_{k})$ and then renormalize:
$\hat{\pi}^{i}_{k}\ :=\ \frac{\tilde{p}(\pi^{i}_{k})}{\sum_{q=1}^{m_{i}}\tilde{p}(\pi^{i}_{q})},\qquad k=1,\dots,m_{i}.$
The perceived lottery under PW is
$\pi_{\textsc{PW}}(L^{i};L^{j})\ :=\ \big(z^{i},\ \hat{\pi}^{i}\big).$
Thus small probabilities are overweighted (pushed up to $0.2$ before renormalization), and large probabilities are underweighted (pulled down to $0.8$). The decision maker only tracks whether an event is “rare,” “intermediate,” or “almost certain” and replaces the exact $p$ by the corresponding internal code.
Context-Dependent Rules
For context-dependent rules we use the state-space representation from Section 2. Given a menu $A_{t}=\{L^{1}_{t},L^{2}_{t}\}$, fix a finite state space
$S=\{1,\dots,n\},\qquad p=(p_{s})_{s\in S}\in\Delta^{n-1},$
and state-contingent payoffs $\big(u^{1}_{t}(s),u^{2}_{t}(s)\big)_{s\in S}$ such that each lottery $L^{i}_{t}$ can be written as the act
$L^{i}_{t}\equiv\big(u^{i}_{t}(s)\big)_{s\in S},\qquad\Pr(s)=p_{s},\quad i\in\{1,2\}.$
As discussed earlier, this is without loss of generality: for any pair of finite-support lotteries one can construct a common finite state space and objective probabilities such that both lotteries are simple functions on that space.
Following the salience literature, comparisons across states are driven by a contrast function. For real numbers $x,y\in\mathbb{R}$, define
c(x,y):=|xy||x|+|y|+1.c(x,y)\ :=\ $\frac${|x-y|}{\,|x|+|y|+1\,}.
This contrast is increasing in the payoff gap |xy||x-y| but exhibits diminishing sensitivity as the absolute magnitudes $|x|$ and $|y|$ grow.
Salience Thinking Rule (SAL).
Salience-based models stress that decision makers overweight states that stand out relative to other states (e.g., bordalo2012_risk; bordalo_et_al2013; salience_review_22). Given $(L^{1},L^{2})$ represented on $(S,p)$, the salience of state $s$ is
$\sigma_{s}(L^{1},L^{2})\ :=\ c\!\big(u^{1}_{t}(s),u^{2}_{t}(s)\big).$
Let $s^{\mathrm{sal}}(L^{1},L^{2})$ be any state attaining the maximal contrast:
$s^{\mathrm{sal}}(L^{1},L^{2})\in\arg\max_{s\in S}\ \sigma_{s}(L^{1},L^{2}),$
with an arbitrary tie-breaking rule. The SAL rule collapses each lottery to its payoff in the most salient state:
$\pi_{\textsc{SAL}}(L^{i};L^{j})\ :=\ \big(u^{i}_{t}(s^{\mathrm{sal}}(L^{1},L^{2}))\ \text{with prob.\ }1\big),\qquad i\in\{1,2\}.$
Under this rule the decision maker focuses entirely on the state with the strongest payoff contrast between $L^{1}$ and $L^{2}$, and then compares lotteries only on their consequences in that state. This is a coarse state-space analogue of salience-based choice à la bordalo2012_risk: instead of reweighting all states based on a salience ranking, we let the most salient state behave “as if” it were the only one perceived.
Regret Rule (REG).
Regret theory emphasizes unfavorable comparisons to forgone outcomes (loomes_sugden86; loomes_sugden1987; loomes92; bleichrodt2015). For each state $s$, the instantaneous regret of choosing $L^{i}$ rather than $L^{j}$ is the foregone payoff difference
$r_{s}^{i}(L^{1},L^{2})\ :=\ \max\{\,u^{j}_{t}(s)-u^{i}_{t}(s),\ 0\,\}.$
We measure the overall severity of potential regret for $L^{i}$ by the worst such payoff loss:
$\rho^{i}(L^{1},L^{2})\ :=\ \max_{s\in S}\ r_{s}^{i}(L^{1},L^{2}).$
Large $\rho^{i}$ means that there exists a state in which the alternative $L^{j}$ delivers a strictly higher payoff than $L^{i}$, and this gap is as large as possible.
The REG rule maps $L^{i}$ to a sure payoff equal to minus this worst regret severity:
$\pi_{\textsc{REG}}(L^{i};L^{j})\ :=\ \big(-\rho^{i}(L^{1},L^{2})\ \text{with prob.\ }1\big).$
Thus lotteries that can be badly beaten by their rival in some state are perceived as having lower certain value. When comparing $(L^{1},L^{2})$, the decision maker behaves as if she were minimizing the maximum foregone payoff difference.
Disappointment Rule (DIS).
Disappointment-based models evaluate outcomes relative to foregone outcomes of the same lottery (e.g., bell85; Gul1991; KoszegiRabin2006; koszegi_rabin2007).
For any lottery $L^{i}$ represented on $(S,p)$ by payoffs $\big(u^{i}_{t}(s)\big)_{s\in S}$, define the reference payoff as the modal payoff of $L^{i}$ (the payoff realized with highest probability). If several payoffs are tied for modal probability, select the highest among the tied values. Denote this reference by $\mathrm{ref}(L^{i})$:
$\mathrm{ref}(L^{i})\ :=\ \max\Big\{\,z:\ \Pr\big(u^{i}_{t}(s)=z\big)=\max_{y}\Pr\big(u^{i}_{t}(s)=y\big)\Big\}.$
Intuitively, $\mathrm{ref}(L^{i})$ captures the outcome the decision maker most expects; when a lower outcome materializes, it triggers disappointment relative to this modal anchor (with the tie-break chosen to maximize the scope for disappointment).
We measure the severity of within-lottery disappointment by the largest contrast from the reference to any strictly lower payoff:
$D(L^{i})\ :=\ \max\Big\{\,c\!\big(\mathrm{ref}(L^{i}),z\big):\ z\in\{u^{i}_{t}(s):s\in S\},\ z<\mathrm{ref}(L^{i})\Big\},$
with the convention $D(L^{i})=0$ if $u^{i}_{t}(s)\geq\mathrm{ref}(L^{i})$ for all $s\in S$ (equivalently, if there is no strictly lower payoff).
The DIS rule maps each alternative to a sure payoff equal to minus this worst reference-loss contrast:
$\pi_{\textsc{DIS}}(L^{i};L^{j})\ :=\ \big(-D(L^{i})\ \text{with prob.\ }1\big).$
Thus, among two lotteries, the one with the smaller maximum disappointment relative to its own modal reference is perceived as better. Because $\mathrm{ref}(L^{i})$ depends only on identifying the modal payoff (with a simple tie-break), the rule avoids detailed probability calculations beyond locating the mode.
Limited Attention Rules
The last group of rules captures extreme forms of limited attention or consideration. [^5] For models of rational inattention and consideration sets, see, among others, sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014; caplin_dean2015. Fix a large constant $M>0$ such that M-M is strictly below the minimum payoff appearing anywhere in the dataset (e.g., $M>\max_{t\in\mathcal{T}}\max_{z\in\text{Supp}(A_{t})}\mid z\mid$).
Attention to $L^{1}$ (A1).
Under Attention to $L^{1}$, the perceived menu replaces the second lottery by a very poor sure outcome:
$\pi_{\textsc{A1}}(L^{1};L^{2})=L^{1},\qquad\pi_{\textsc{A1}}(L^{2};L^{1})=\big(-M\ \text{with prob.\ }1\big).$
Thus $L^{1}$ is evaluated as given, while $L^{2}$ is collapsed to a dominated baseline. Intuitively, the decision maker pays attention only to the first alternative in the menu and disregards the second.
Attention to $L^{2}$ (A2).
Symmetrically, under Attention to $L^{2}$ the first lottery is replaced by the dominated baseline, while $L^{2}$ is perceived without distortion:
[…]
These two attention rules do not correspond to a specific structural model of risk preferences; instead, they serve as coarse “default” heuristics encoding extreme one-sided attention. In our implementation, we always include both A1 and A2 in the library.
Summary. Our baseline library is
$\mathcal{F}=\{\textsc{Id},\textsc{MMn},\textsc{MMx},\textsc{MAP},\textsc{PW},\textsc{SAL},\textsc{REG},\textsc{DIS},\textsc{A1},\textsc{A2}\}.$
The library is deliberately simple and nonparametric: all ten rules are parameter-free and are designed as coarse, reduced-form implementations of four broad mechanisms central in the bounded-rationality and behavioral-decision literatures. MMn, MMx and MAP implement outcome-simplification heuristics (rubinstein88; gigerenzer_gaissmaier2011; brandstatter_et_al2006) by projecting each lottery onto a single payoff attribute (minimum, maximum, or modal payoff). PW distorts probabilities in a coarse fashion that echoes probability weighting in cumulative prospect theory (kahneman_tversky1979; tversky_kahneman1992; prelec1998). SAL, REG and DIS capture different forms of context-dependent comparisons across and within lotteries, providing nonparametric counterparts to salience, regret and disappointment models (bordalo2012_risk; LoomesSugden1986; bell85; Gul1991; KoszegiRabin2006). Finally, A1 and A2 encode extreme one-sided attention and thus serve as simple menu-level representations of limited attention and consideration-set formation (sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014; caplin_dean2015). All ten rules fit the general menu-transformer framework $f(L^{1},L^{2})=(\pi_{f}(L^{1};L^{2}),\pi_{f}(L^{2};L^{1}))$. The rule library introduced in this section is tailored to binary lottery choice environments. There might be additional relevant rules for more complex choice environments, including sampling-based perception, support compression, or rank-dependent distortions beyond a single probability. [^6] See, for example, hu2023. The next subsection specifies when a profile of such rules is considered an admissible explanation of a dataset, and how we quantify the parsimony of these explanations.
2.2 Identification
[…]
$\mathbf{f}=(f_{t})_{t\in\mathcal{T}}\ \in\ \mathcal{F}^{\mathcal{T}}$
assigns one rule to each observed menu. At observation $t\in\mathcal{T}$, rule $f_{t}$ transforms the objective menu $A_{t}=\{L_{t}^{1},L_{t}^{2}\}$ into a perceived menu
$\widetilde{A}_{t}(\mathbf{f})=\{\widetilde{L}^{1}_{t,f_{t}},\widetilde{L}^{2}_{t,f_{t}}\},\qquad\widetilde{L}^{i}_{t,f}:=\pi_{f}(L_{t}^{i};L_{t}^{-i}),\quad i\in\{1,2\},$
where $L_{t}^{-i}$ denotes the alternative lottery in menu $A_{t}$ (i.e., $L_{t}^{-1}=L_{t}^{2}$ and $L_{t}^{-2}=L_{t}^{1}$). The realized choice $x_{t}=L_{t}^{i_{t}}$ is perceived as
$\widetilde{x}_{t}(\mathbf{f}):=\pi_{f_{t}}(x_{t};L_{t}^{-i_{t}}).$
The analyst observes only $D=\{(x_{t},A_{t})\}_{t\in\mathcal{T}}$, while the decision maker behaves as if choosing $\widetilde{x}_{t}(\mathbf{f})\in\widetilde{A}_{t}(\mathbf{f})$ in each perceived problem. We denote by
$\widetilde{D}(\mathbf{f}):=\big\{(\widetilde{x}_{t}(\mathbf{f}),\widetilde{A}_{t}(\mathbf{f}))\big\}_{t\in\mathcal{T}}$
the perceived dataset generated by $\mathbf{f}$. Our identification strategy specifies which rule profiles $\mathbf{f}$ qualify as acceptable explanations of $D$.
Local discriminability.
[…]
$\widetilde{A}_{t}(f)\ :=\ \big\{\pi_{f}(L_{t}^{1};L_{t}^{2}),\,\pi_{f}(L_{t}^{2};L_{t}^{1})\big\},\qquad\widetilde{x}_{t}(f)\ :=\ \pi_{f}(x_{t};L_{t}^{-i_{t}}).$
Let $\widetilde{L}_{t}^{-}(f)$ denote the perceived alternative to the realized choice under $f$:
$\widetilde{L}_{t}^{-}(f)\ :=\ \pi_{f}(L_{t}^{-i_{t}};x_{t}).$
We define the set of strictly discriminating rules at $t$ as
$\mathcal{F}^{\mathrm{strict}}_{t}\ :=\ \Big\{\,f\in\mathcal{F}:\ \widetilde{x}_{t}(f)\ \triangleright\ \widetilde{L}_{t}^{-}(f)\,\Big\}.$
[…]
Assumption 1 (Local Discriminability).
For each observation $t\in\mathcal{T}$, the rule used at $t$ belongs to the strictly discriminating set, $f_{t}\in\mathcal{F}^{\mathrm{strict}}_{t}$.
Assumption 1 formalizes the idea that rules provide a “clear-cut line of conduct” at the menu level: whenever it is possible to interpret a choice as selecting a strictly dominant perceived option under some rule, only such rules are admissible. Importantly, this does not require dominance in the objective menu; dominance is evaluated after the rule-induced transformation. The requirement is therefore entirely about the internal coherence of the rule-based description. Because the attention rules A1 and A2 are always in $\mathcal{F}$, the set $\mathcal{F}^{\mathrm{strict}}_{t}$ is never empty: at any binary menu, at least one of the two attention rules makes the chosen alternative strictly dominate the other in the perceived menu. This yields the following simple completeness property.
Proposition 1.
For any $t\in\mathcal{T}$, we have $\mathcal{F}^{\mathrm{strict}}_{t}\neq\varnothing$.
In Appendix A, we show that if the rule profile $\mathbf{f}$ satisfies Assumption 1, then the perceived dataset satisfies a cyclical consistency condition that generalizes the standard SARP (Strong Axiom of Revealed Preference) to our choice environment. This implication marks a fundamental difference with the standard revealed-preference approach. In classical utility-based models, cyclical consistency is imposed directly on the observed dataset $D$ and serves as a common weak denominator across broad families of theories, including expected utility theory and its generalizations. The data are required to look as if they were generated by a stable underlying ordering. In our framework, the raw dataset $D$ is not expected to satisfy such consistency with respect to the normative benchmark. Indeed, many canonical “anomalies” correspond precisely to monotonicity violations or to revealed-preference cycles under FSD. The discipline instead comes from a different place. Behavior is generated by menu-dependent rules that transform each problem into a perceived one, and the observed choice must be locally defensible as a strict FSD improvement in that perceived problem. Under FSD, this local discipline is strong enough to guarantee that the induced perceived dataset $\widetilde{D}(\mathbf{f})$ satisfies a strong cyclical consistency condition. The empirical content therefore shifts away from asking whether choices are globally consistent as observed, and toward asking whether they can be organized by a parsimonious collection of simple transformations that make them locally dominance-improving once perception is taken into account.
3 Examples
This section shows how familiar “anomalies” in risky choice can be viewed through our library. Each example illustrates which rules in $\mathcal{F}$ satisfy Local Discriminability.
3.1 FSD violation
Consider a single choice between lotteries $L^{1}$ and $L^{2}$ defined on three equally likely states $s\in\{1,2,3\}$. Lottery $L^{2}$ strictly first-order stochastically dominates $L^{1}$ for any $\varepsilon>0$: it yields a (weakly) higher payoff in every state and a strictly higher payoff in at least one state. Choosing $L^{1}$ therefore violates monotonicity with respect to $\geq_{\mathrm{FSD}}$.
Table 1: Violation of FSD and regret/salience rules
state $s$ 11 22 33
$p_{s}$ $1/3$ $1/3$ $1/3$
$L^{1}$ 2020 0 $10$
$L^{2}$ 0 $10+\varepsilon$ 2020
Empirically, such choices are not rare; see, e.g., bleichrodt2015. A standard explanation is regret. Under our regret rule REG, the severity of regret from choosing $L^{i}$ against $L^{j}$ is the largest state-wise payoff loss
$\rho^{i}(L^{1},L^{2})\ :=\ \max_{s\in S}\max\{\,z^{j}_{s}-z^{i}_{s},\,0\,\},$
where $(z^{1}_{s},z^{2}_{s})_{s\in S}$ denotes a common state-space representation of $(L^{1},L^{2})$ as in Section 2. In the menu $\{L^{1},L^{2}\}$, choosing $L^{2}$ exposes the decision maker to a potential forgone gain of $20$ in state $s=1$, whereas choosing $L^{1}$ exposes her only to a forgone gain of $10+\varepsilon$ in state $s=2$. Hence $\rho^{2}(L^{1},L^{2})>\rho^{1}(L^{1},L^{2})$ and REG strictly prefers $L^{1}$ to $L^{2}$.
The salience rule SAL delivers a distinct but related explanation. It computes the state-wise contrasts $c(z^{1}_{s},z^{2}_{s})$ and focuses attention on the most salient state. Here the largest payoff gap is again in state $s=1$ (2020 vs. 0), so SAL collapses the comparison to a sure $20$ under $L^{1}$ versus a sure 0 under $L^{2}$ and strictly prefers $L^{1}$. In terms of our library, this FSD violation can be explained by REG or SAL, while the other rules fail to yield a strict ranking (e.g., MMn yields 0 for both lotteries; MMx yields $20$ for both; DIS delivers the same disappointment severity for both; and MAP and PW do not discriminate).
3.2 Cyclical consistency (I)
Our next example, from bleichrodt2015 following loomes_sugden1987 and sugden1991, illustrates how behavior consistent with regret and salience can generate a revealed-preference cycle under $\geq_{\mathrm{FSD}}$ even though no FSD violation occurs at any single menu. There are four equiprobable states and four lotteries $L^{3},\dots,L^{6}$:
Table 2: Preference cycle (I)
state $s$ 11 22 33 44
$p_{s}$ $1/4$ $1/4$ $1/4$ $1/4$
$L^{3}$ 0 $10$ 2020 $30$
$L^{4}$ $30$ 0 $10$ 2020
$L^{5}$ 2020 $30$ 0 $10$
$L^{6}$ $10$ 2020 $30$ 0
The decision maker faces the menus
$\mathcal{A}=\big\{\{L^{3},L^{4}\},\ \{L^{4},L^{5}\},\ \{L^{5},L^{6}\},\ \{L^{6},L^{3}\}\big\},$
and we observe
$x_{1}=L^{4},\quad x_{2}=L^{5},\quad x_{3}=L^{6},\quad x_{4}=L^{3}.$
This is a classical example of cyclical consistency violation. [^7] With the notations of Appendix A, under $\geq_{\mathrm{FSD}}$, all four lotteries are pairwise incomparable: they share the same multiset of payoffs $\{0,10,20,30\}$ with equal probabilities. No single choice therefore violates FSD, and the strict revealed-preference relation $P$ (defined from $\geq_{\mathrm{FSD}}$) is empty. However, the weak relation $R$ from Section 2 satisfies $L^{4}\,R\,L^{3},\quad L^{5}\,R\,L^{4},\quad L^{6}\,R\,L^{5},\quad L^{3}\,R\,L^{6},$ so the objective dataset induces a nontrivial revealed-preference cycle (and hence fails the SARP-type condition in Definition 10, Appendix A). The observed choices are exactly reproduced by REG and SAL. Under REG, in each menu $\{L^{i},L^{j}\}$ the rejected lottery has regret severity $30$ while the chosen one has regret severity $10$. For instance, in $\{L^{3},L^{4}\}$ the regret from choosing $L^{3}$ is $30$ (state $1$), whereas the regret from choosing $L^{4}$ is $10$ (states 22-44). SAL behaves similarly: in each pair the most salient state is the one where one lottery yields $30$ and the other 0, and SAL chooses the $30$-paying lottery in that state, again generating the 4-cycle. By contrast, MMn, MMx, DIS, MAP, and PW do not strictly discriminate in these menus.
3.3 Common-consequence Allais paradox
We next connect the Allais common-consequence pattern to our rule library. Consider the standard pair of choice problems. In the first problem, subjects choose between
[…]
and in the second problem between
[…]
In this well-known example, many subjects choose $L^{7}$ over $L^{8}$ and $L^{10}$ over $L^{9}$, violating the independence axiom of expected utility. Within our rule library, SAL and REG provide natural explanations. A salience thinker focuses on the possibility of winning \$2,500 versus \$0 in the first problem, and on the possibility of receiving \$0 versus \$2,400 in the second. A regret-based thinker compares worst forgone outcomes across the two menus. In particular, under REG, in the first choice problem the decision maker prefers $L^{7}$ because choosing $L^{8}$ entails a worst forgone gain of \$2,500 (relative to the state in which $L^{7}$ pays \$2,500 and $L^{8}$ pays \$0). In the second choice problem, the sure \$2,400 outcome has strictly lower regret severity.
3.4 Common-ratio Allais paradox
Consider now the common-ratio pair. In the first problem, subjects choose between
[…]
and in the second problem between
[…]
The classical pattern is $L^{11}$ chosen over $L^{12}$ and $L^{13}$ chosen over $L^{14}$. No single rule from $\mathcal{F}$ can explain both menus while satisfying Local Discriminability. PW, for instance, is often invoked for common-ratio patterns, but here its coarse coding of probabilities does not yield a strict ranking in the first problem. The modal-payoff rule MAP, on the other hand, does. In $\{L^{11},L^{12}\}$ the modal payoff for $L^{11}$ is \$3,000 (probability $0.95$), while the modal payoff for $L^{12}$ is \$0 (probability $0.525$), so MAP strictly prefers $L^{11}$. In the second menu, SAL and REG both favor $L^{13}$: they focus on the contrast between \$6,000 and \$0, or on the worst forgone loss from not choosing $L^{13}$. Accordingly, a locally discriminating rationalization uses MAP in the first problem and SAL or REG in the second.
3.5 The reflection effect
We now turn to the reflection effect: risk aversion over fair gains and risk seeking over fair losses. Consider two 50-50 choices over gains and losses. In the gain domain:
$L^{15}=\begin{cases}100&\text{with prob. }1,\end{cases}\qquad L^{16}=\begin{cases}200&\text{with prob. }0.5,\\ 0&\text{with prob. }0.5,\end{cases}$
and in the loss domain:
$L^{17}=\begin{cases}-100&\text{with prob. }1,\end{cases}\qquad L^{18}=\begin{cases}0&\text{with prob. }0.5,\\ -200&\text{with prob. }0.5.\end{cases}$
All gambles are fair: $\mathbb{E}[L^{15}]=\mathbb{E}[L^{16}]=100$ and $\mathbb{E}[L^{17}]=\mathbb{E}[L^{18}]=-100$. A common pattern selects $L^{15}$ and $L^{18}$.
In our library, REG predicts indifference in both problems: regret severity is the same for the two lotteries in each menu, so REG cannot generate the reflection effect. SAL, by contrast, is consistent with the observed pattern. In $\{L^{15},L^{16}\}$, the largest contrast is between receiving \$100 under $L^{15}$ and \$0 under $L^{16}$, leading to risk aversion. In $\{L^{17},L^{18}\}$, the most salient gap is between a loss of \$100 and no loss (0), pushing choice toward the risky $L^{18}$. The key driver is diminishing sensitivity in the contrast function: the gap between 0 and $\pm 100$ has higher contrast than that between $\pm 100$ and $\pm 200$.
3.6 Cyclical consistency (II)
Our next example shows a pure salience-driven violation of strong cyclical consistency. Consider three equiprobable states and three lotteries $L^{19},L^{20},L^{21}$:
Table 3: Three-lottery salience cycle
state $s$ 11 22 33
$p_{s}$ $1/3$ $1/3$ $1/3$
$L^{19}$ 0 0 2020
$L^{20}$ $10$ 88 0
$L^{21}$ 0 $30$ 0
The menus are
$\mathcal{A}=\big\{\{L^{19},L^{20}\},\ \{L^{20},L^{21}\},\ \{L^{21},L^{19}\}\big\},$
and suppose the choices are $L^{19}$, $L^{20}$, and $L^{21}$ in respectively the first, second, and third choice problem. [^8] Since $L^{19}\neq L^{21}$, the dataset violates the strong cyclical consistency condition of Appendix A. Under SAL, in $\{L^{19},L^{20}\}$ the most salient gap is in $s=3$ (2020 vs. 0), so $L^{19}$ is chosen. In $\{L^{20},L^{21}\}$, the contrast in $s=1$ is $c(10,0)=10/11$, larger than the contrast in $s=2$, $c(8,30)=22/39$, so $L^{20}$ is chosen. In $\{L^{21},L^{19}\}$, the largest gap is in $s=2$ ($30$ vs. 0), so $L^{21}$ is chosen. REG does not reproduce the same cycle because regret severity for $L^{20}$ in $\{L^{20},L^{21}\}$ comes from state $2$ (foregone $30$), which exceeds the regret severity for $L^{21}$; REG therefore prefers $L^{21}$ to $L^{20}$ in that menu. MMn, MMx, and DIS also treat the three lotteries more symmetrically and do not yield a 3-cycle. This example isolates a pattern that is salience-based.
3.7 Mixture reversal
Finally, we consider a mixture reversal: a preference reversal induced by mixing both options with a common baseline. In the first problem, the decision maker chooses between a safe and a risky gain:
$L^{24}=\begin{cases}200&\text{with prob. }1\end{cases}\qquad\text{vs}\qquad L^{25}=\begin{cases}300&\text{with prob. }0.8\\ 0&\text{with prob. }0.2\;.\end{cases}$
In the second problem, both lotteries are mixed with a sure-zero baseline $L^{0}=(0\ \text{with prob. }1)$ with weight $\lambda=0.6$, $L^{26}=\lambda L^{24}+(1-\lambda)L^{0}$ and $L^{27}=\lambda L^{25}+(1-\lambda)L^{0}$:
[…]
Suppose that we observe the choices $L^{24}$ and $L^{27}$, the standard mixture reversal pattern.
Within our library, DIS gives a natural explanation. In the first problem, $L^{24}$ has modal payoff $ref(L^{24})=200$ and no lower outcome, so $D(L^{24})=0$. Lottery $L^{25}$ has modal payoff $ref(L^{25})=300$ and a lower outcome 0, so $D(L^{25})=c(300,0)>0$, and DIS prefers the sure outcome. In the second problem, $L^{26}$ has modal outcome $200$ and lower outcome 0, so $D(L^{26})=c(200,0)>0$, whereas $L^{27}$ has modal outcome 0 (probability $0.52$) and no strictly lower outcome, so $D(L^{27})=0$. DIS therefore reverses its ranking and now prefers $L^{27}$. The mixture with $L^{0}$ flips the modal (reference) payoff for the risky option but not for the safe one, eliminating disappointment for $L^{27}$ and introducing it for $L^{26}$.
SAL and REG can also rationalize the reversal. In the first problem, both focus on the event where $L^{25}$ yields 0 and $L^{24}$ yields $200$, favoring the safe option; in the second, they focus on the event where $L^{27}$ yields $300$ while $L^{26}$ yields 0, favoring the risky option. By contrast, MMn selects the safe option in the first problem but is indifferent in the second (both minima are zero), MMx always prefers the risky option, and MAP predicts $L^{25}$ in the first problem and $L^{26}$ in the second.
Discussion
Each example isolates which rules in the library $\mathcal{F}$ can explain the observed choices while satisfying Local Discriminability. Some patterns are compatible with several distinct rules (e.g., the common-consequence pair can be explained by either salience or regret), while others are not. For instance, the reflection effect separates salience from regret, since REG predicts indifference whereas SAL delivers the observed switch from risk aversion over gains to risk seeking over losses; the common-ratio pair forces the use of the modal-payoff rule MAP in the first menu and rules such as SAL or REG in the second; and the three-lottery cycle (Cyclical consistency II) is essentially salience-specific. Whenever such patterns are present in the data, they mechanically restrict which rules can be used at the corresponding menus in any admissible profile.
Taken together, the examples show that a small, nonparametric library of rules can organize a broad range of well-known behaviors in risky choice, and that different patterns load on different subsets of rules. In the next section, we move from these qualitative illustrations to a quantitative complexity measure that asks how parsimoniously a whole dataset explained by such rules, and how much of this parsimony is carried by specific rules.
4 Maximum Rule Concentration Index (MRCI)
We now formalize how parsimoniously a dataset can be explained by our rule library.
4.1 Admissible profiles and rule assignments
Let the dataset be
$D=\{(x_{t},A_{t})\}_{t\in\mathcal{T}},\qquad A_{t}=\{L_{t}^{1},L_{t}^{2}\},\qquad T:=|\mathcal{T}|.$
A rule profile $\mathbf{f}=(f_{t})_{t\in\mathcal{T}}\in\mathcal{F}^{\mathcal{T}}$ induces the perceived dataset
$\widetilde{D}(\mathbf{f})\ :=\ \Big\{\big(\widetilde{x}_{t}(\mathbf{f}),\widetilde{A}_{t}(\mathbf{f})\big)\Big\}_{t\in\mathcal{T}},$
where
$\widetilde{A}_{t}(\mathbf{f})=\{\widetilde{L}^{1}_{t,f_{t}},\widetilde{L}^{2}_{t,f_{t}}\},\qquad\widetilde{L}^{i}_{t,f}:=\pi_{f}(L_{t}^{i};L_{t}^{-i}),\qquad\widetilde{x}_{t}(\mathbf{f}):=\pi_{f_{t}}(x_{t};L_{t}^{-i_{t}}).$
As in Assumption 1, an admissible profile must, at each observation $t\in\mathcal{T}$, use a rule that strictly favors the chosen option in the perceived menu. It is convenient to work with rule assignments. For each $t\in\mathcal{T}$ and $f\in\mathcal{F}$, let $y_{t,f}\in\{0,1\}$ indicate whether rule $f$ is used at observation $t$. An assignment is a matrix $\mathbf{y}=\{y_{t,f}\}_{t\in\mathcal{T},\,f\in\mathcal{F}}$ satisfying
$\sum_{f\in\mathcal{F}}y_{t,f}=1,\qquad t\in\mathcal{T}.$
Each assignment $\mathbf{y}$ corresponds one-to-one to a profile $\mathbf{f}$ on $\mathcal{T}$.
Definition 2 (Admissible assignment).
Let $\mathcal{C}(D;\geq_{FSD},\mathcal{F})$ be the set of rule assignments $\mathbf{y}=\{y_{t,f}\}$ defined on $\mathcal{T}$ such that, for each $t\in\mathcal{T}$, exactly one rule is used and that rule is strictly discriminating at $t$:
$\sum_{f\in\mathcal{F}}y_{t,f}=1,\qquad\sum_{f\in\mathcal{F}_{t}^{\mathrm{strict}}}y_{t,f}=1,\qquad t\in\mathcal{T}.$
Elements of $\mathcal{C}(D;\geq_{FSD},\mathcal{F})$ are called admissible rule assignments.
Definition 2 restates our identifying restriction (Local Discriminability) in terms of assignment variables. Henceforth, we write $\mathcal{C}(D)$ as shorthand when feasible.
4.2 The Maximum Rule Concentration Index
Given an admissible assignment $\mathbf{y}\in\mathcal{C}(D)$, define the frequency with which rule $f$ is used as
$s_{f}(\mathbf{y}):=\frac{1}{T}\sum_{t\in\mathcal{T}}y_{t,f},\qquad f\in\mathcal{F},$
and the associated Herfindahl-Hirschman index (HHI) by
$\mathrm{HHI}(\mathbf{y}):=\sum_{f\in\mathcal{F}}s_{f}(\mathbf{y})^{2}\ \in\ \Big[\tfrac{1}{\mid\mathcal{F}\mid},1\Big].$
$\mathrm{HHI}(\mathbf{y})$ equals $1/\mid\mathcal{F}\mid$ when all rules are used equally often, and equals $1$ when a single rule accounts for all observations in $\mathcal{T}$.
Definition 3 (Maximum Rule Concentration Index).
The Maximum Rule Concentration Index is
$\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})\ :=\ \max_{\mathbf{y}\in\mathcal{C}(D;\geq_{FSD},\mathcal{F})}\mathrm{HHI}(\mathbf{y}),$
and the effective number of rules is $N_{\mathrm{eff}}:=1/\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})$.
The MRCI measures how concentrated the explanation of $D$ can be. If $\operatorname{MRCI}$ is close to $1$, a single heuristic dominates; if $\operatorname{MRCI}$ is low, admissible explanations necessarily spread weight across many rules. Let
$\alpha\ :=\ \frac{1}{T}\sum_{t\in\mathcal{T}}\mathbf{1}\{x_{t}=L_{t}^{1}\}$
denote the share of observations in which the decision maker chooses the first-listed lottery in the menu.
Basic properties.
Proposition 2.
For any dataset $D$, and library $\mathcal{F}$:
  1. i
    Since $\{\textsc{A1},\textsc{A2}\}\subseteq\mathcal{F}$,
    $\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})\in\big[\alpha^{2}+(1-\alpha)^{2},\,1\big].$
  2. $i$
    $\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})=1$ if and only if there exists a rule $f\in\mathcal{F}$ such that
    $f\in\mathcal{F}_{t}^{\mathrm{strict}}\qquad\text{for all }t\in\mathcal{T}.$
  3. iii
    Concatenating $k$ identical copies of $D$ (i.e., repeating each observation $k$ times) leaves $\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})$ unchanged.
  4. iv
    If for some $t\in\mathcal{T}$ there is a unique strictly discriminating rule $f$ (i.e., $\mathcal{F}_{t}^{\mathrm{strict}}=\{f\}$), then every maximizer $\mathbf{y}^{\star}$ of $\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})$ satisfies $y^{\star}_{t,f}=1$.
The upper bound in (i) corresponds to a one-rule explanation. The lower bound comes from the always-feasible “attention” assignment: use A1 whenever $x_{t}=L_{t}^{1}$ and A2 otherwise, which yields $\mathrm{HHI}=\alpha^{2}+(1-\alpha)^{2}$. Property (iii) shows that MRCI is intensive (composition-based) rather than extensive (sample-size-based). Property (iv) highlights “forced moves” that anchor the optimization: if an observation can only be strictly discriminated by a single rule, every admissible maximizer must assign that observation to the same rule.
Monotonicity
Let $\mathcal{F}^{\prime}\supseteq\mathcal{F}$ be an expanded library.
Proposition 3 (Monotonicity).
For any dataset $D$,
$\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F}^{\prime})\ \geq\ \operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F}).$
Moreover, a necessary condition for strict inequality is that there exists an MRCI-maximizing assignment $\mathbf{y}^{\star}$ under $\mathcal{F}$ and a new rule $f\in\mathcal{F}^{\prime}\setminus\mathcal{F}$ such that, on a nonempty subset of observations, $f$ is strictly discriminating and can replace (while maintaining admissibility) two distinct rules $r\neq s$ that both have positive shares under $\mathbf{y}^{\star}$ (i.e., $s_{r}(\mathbf{y}^{\star}),s_{s}(\mathbf{y}^{\star})>0$).
Intuitively, a new rule increases concentration only if it can act as a “unifier,” substituting for multiple distinct positive-share rules on the observations where those rules were previously needed.
Stability Score
Because admissible assignments and MRCI–maximizers need not be unique, we look at how much concentration can be preserved when rules are deleted. Fix a finite family $\mathcal{K}$ of permutations of $\mathcal{F}$. For each order $k\in\mathcal{K}$, we construct a Deletion–Stable Set $E^{(k)}$ by iteratively removing rules $f^{(k)}_{i}$ if the MRCI on the restricted library remains equal to the original optimum $H^{\star}=\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})$. The final set $E^{(k)}$ is a minimal library that maintains the coherence of the originally explained behavior.
Definition 4 (Stability Score).
For any rule $f\in\mathcal{F}\setminus\{\textsc{A1},\textsc{A2}\}$, the Stability Score of $f$ is
$\kappa(f):=\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}\mathbf{1}\{f\in E^{(k)}\}\in[0,1].$
The score $\kappa(f)$ measures how often $f$ is indispensable for attaining $H^{\star}$. If $\kappa(f)=1$, removing $f$ always breaks the optimal concentration structure; $f$ is a pillar of the subject’s rule usage. If $\kappa(f)=0$, the concentration provided by $f$ can always be replicated by other rules in the library.
Concentration Gain
The Stability Score measures indispensability; the Concentration Gain measures the marginal contribution to coherence.
Definition 5 (Concentration Gain).
For any rule $f\in\mathcal{F}\setminus\{\textsc{A1},\textsc{A2}\}$, the Concentration Gain of $f$ is
$\phi(f)\ :=\ \frac{\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})\;-\;\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F}\setminus\{f\})}{\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})}\in[0,1].$
A value $\phi(f)=0$ means that removing $f$ does not affect the coherence of the explanation. A high $\phi(f)$ indicates that $f$ is a "unifying" rule: without it, the subject’s behavior fragments into a collection of disparate heuristics (or becomes largely unexplainable), causing the concentration index to collapse. Together, the MRCI (Coherence), $\kappa(f)$ (Robustness), and $\phi(f)$ (Impact) provide a complete decomposition of the decision maker’s strategy.
4.3 Examples, revisited
To illustrate these notions, it is useful to revisit the examples from Section 3. For each pattern we can identify (i) which rules are locally admissible at each menu, (ii) whether a one-rule explanation exists, and (iii) which rules become indispensable once we pool examples.
Single-example datasets.
For most canonical patterns taken in isolation (e.g., the FSD violation, the two cyclical examples, and the common-consequence pair), the admissible set $\mathcal{C}(D)$ contains assignments that use only one non-attention rule from $\mathcal{F}$ across all observations. For instance, in the FSD-violation example, either REG or SAL can be used at the single menu in a way that satisfies Local Discriminability. For such small datasets, MRCI therefore cannot distinguish sharply between candidate rules: several heuristics are perfect substitutes for explaining the pattern under Local Discriminability, which leads to $\phi(f)=0$ for each of them and stability scores that spread symmetrically across the admissible rules. Isolated anomalies mainly tell us that some simple rule can account for the behavior while preserving perceived rationality; they do not yet pin down which rule is responsible.
A pooled illustration.
The value of MRCI and the rule-usage diagnostics becomes more apparent when we pool multiple examples, as in standard datasets. Let $D_{\mathrm{pool}}$ collect the sixteen observations spanning: the FSD violation; the four-lottery cycle (Cyclical consistency I); the common-consequence Allais pair; the common-ratio pair; the reflection effect (two menus); the three-lottery salience cycle (Cyclical consistency II); and the mixture reversal (two menus). For each observation, Table 4 summarizes which non-attention rules in $\mathcal{F}$ are locally admissible (i.e., belong to $\mathcal{F}_{t}^{\mathrm{strict}}$) once Local Discriminability is imposed.
Table 4: Summary: examples and locally admissible non-attention rules
Example Admissible rules in $\mathcal{F}_{t}^{\mathrm{strict}}\setminus\{\textsc{A1},\textsc{A2}\}$
FSD violation REG, SAL
Cyclical consistency (I) REG, SAL
Common-consequence REG, SAL
Common-ratio $\{L^{11},L^{12}\}$: MAP; $\{L^{13},L^{14}\}$: REG, SAL
Reflection effect SAL
Cyclical consistency (II) SAL
Mixture reversal DIS, REG, SAL
In this pooled dataset $D_{\mathrm{pool}}$, SAL is locally admissible in fifteen out of sixteen observations; the only exception is the first common-ratio menu $\{L^{11},L^{12}\}$, where only MAP and one attention rule are locally discriminating. There therefore exists an admissible assignment that uses SAL at all observations where it is admissible and MAP at the remaining one. Thus every maximizer $\mathbf{y}^{\star}$ of $\mathrm{HHI}$ satisfies
$\text{HHI}(\mathbf{y}^{\star})=\operatorname{MRCI}(D_{\mathrm{pool}};\geq_{FSD},\mathcal{F})=\Big(\tfrac{15}{16}\Big)^{2}+\Big(\tfrac{1}{16}\Big)^{2}=\tfrac{113}{128}\approx 0.883,$
so that $N_{\mathrm{eff}}\approx 1.13$. The pooled behavior is thus close to being explained by a single heuristic (salience), with MAP or an attention rule playing a purely residual role in one menu.
Stability scores and concentration gains refine this picture. Because attention rules can always rationalize any single observation, MAP is never indispensable for sustaining $H^{\star}=\operatorname{MRCI}(D_{\mathrm{pool}};\geq_{FSD},\mathcal{F})$: there are deletion orders in which MAP can be removed without changing the optimal MRCI value (the corresponding observation can be reassigned to A1 or A2). Hence $\kappa(\textsc{MAP})=0$, and in this pooled illustration $\phi(\textsc{MAP})=0$ as well.
For SAL, removing it makes maximal concentration fall. Indeed, when SAL is excluded from the library, many observations must be reassigned to REG and to the attention rules, which reduces the maximal HHI sharply. A conservative reassignment yields
$\mathrm{HHI}\approx\Big(\tfrac{10}{16}\Big)^{2}+\Big(\tfrac{6}{16}\Big)^{2}\approx 0.53,$
so the implied concentration gain of SAL is at least
$\phi(\textsc{SAL})\ \geq\ \frac{0.883-0.53}{0.883}\ \approx\ 0.4.$
Moreover, SAL appears in every deletion-stable set across permutations, so $\kappa(\textsc{SAL})=1$: it is order-robustly indispensable for sustaining the maximally concentrated explanation.
In this pooled illustration, MRCI indicates a highly parsimonious structure ($N_{\mathrm{eff}}\approx 1$), while the stability score and concentration gain reveal which rule carries this parsimony. Salience emerges as the workhorse decision rule: it is always retained in value-preserving prunings of the library and accounts for a substantial share of the attainable concentration, whereas other rules such as MAP are weakly identified from the standpoint of rule concentration.
5 Computational Strategy
Computing the MRCI amounts to solving a combinatorial problem: we must select, for each menu $t$, a rule $f_{t}\in\mathcal{F}_{t}^{\mathrm{strict}}$ so as to maximize the Herfindahl–Hirschman concentration index
$\mathrm{HHI}(\mathbf{y})=\sum_{f\in\mathcal{F}}s_{f}(\mathbf{y})^{2},\qquad s_{f}(\mathbf{y}):=\frac{1}{T}\sum_{t=1}^{T}y_{t,f},$
subject to the requirement that the assigned rule at each menu satisfies local discriminability. Here $\mathbf{y}=(y_{t,f})_{t,f}$ is the assignment matrix with $y_{t,f}\in\{0,1\}$ and $\sum_{f}y_{t,f}=1$ for all $t$, and the constraint requires that if $y_{t,f}=1$, then $f\in\mathcal{F}_{t}^{\mathrm{strict}}$ (Assumption 1).
An exact formulation of this problem as a mixed–integer quadratic program (MIQP) is provided in Appendix D. However, maximizing a convex objective function (concentration) over discrete assignments is an NP-hard problem. As demonstrated in our benchmarking exercise (see Appendix D), the MIQP approach scales fast with sample size, rendering it computationally intractable for standard datasets. To address this, we employ a heuristic algorithm that exploits the structure of the objective. Since we are maximizing concentration, optimal solutions by convexity tend to load on rules that can rationalize many menus simultaneously. The algorithm below is designed to efficiently explore such high–coverage, high–concentration assignments.
5.1 Overview of the heuristic
The procedure has three components:
  1. i
    a pre–processing step that computes, for each $(t,f)$, whether rule $f$ is locally admissible at menu $t$;
  2. $i$
    a greedy assignment step that, for a given ordering of rules, assigns rules to as many unassigned admissible menus as possible;
  3. iii
    a random–restart step that repeats the greedy pass under different rule orderings, and retains the assignment with the highest $\mathrm{HHI}$.
We now describe each component in turn.
Phase 1: Pre–processing
Local admissibility.
For every menu $t$ and rule $f\in\mathcal{F}$, we compute whether $f$ is strictly discriminating at $t$ in the sense of Assumption 1. Formally, we set
$a_{t,f}:=\mathbf{1}\{f\in\mathcal{F}_{t}^{\mathrm{strict}}\},$
where $\mathcal{F}_{t}^{\mathrm{strict}}$ is the set of rules such that, once applied to $A_{t}$, the perceived chosen lottery $\widetilde{x}_{t}(f)$ strictly dominates its perceived alternative. In all subsequent steps we only consider assignments $(t,f)$ with $a_{t,f}=1$. This step is done once per dataset.
Phase 2: Greedy assignment for a fixed rule order
The core of the heuristic is a greedy assignment procedure that, given an ordering of the rules, attempts to assign each rule to as many admissible menus as possible.
Rule ordering.
Let $(f_{1},\dots,f_{F})$ be an ordering of the rules in $\mathcal{F}$. In the first run, we use a deterministic “popularity–first” order: rules are sorted in decreasing order of
$\text{coverage}(f):=\sum_{t=1}^{T}a_{t,f},$
i.e. the number of menus at which they are strictly discriminating. In subsequent runs (Phase 3) we use random permutations of $\mathcal{F}$ to explore alternative local optima.
State variables.
We maintain an assignment matrix $\mathbf{y}$, initialized at $y_{t,f}=0$ for all $t,f$.
Greedy pass.
We loop over rules in the chosen order $f_{1},\dots,f_{F}$. For a given rule $f_{k}$:
  1. i
    We collect the set of menus where $f_{k}$ is admissible and not yet assigned:
    $\mathcal{T}_{k}:=\big\{t:\ a_{t,f_{k}}=1,\ \sum_{h\in\mathcal{F}}y_{t,h}=0\big\}.$
  2. $i$
    We assign rule $f_{k}$ to all such menus. That is, for all $t\in\mathcal{T}_{k}$, we set $y_{t,f_{k}}=1$.
At the end of this pass, all menus are assigned, as from Proposition 1, the set $\mathcal{F}_{t}^{\mathrm{strict}}$ is never empty. The outcome of Phase 2 is a feasible assignment matrix $\mathbf{y}^{(k)}$ for the given rule order, along with its value $\mathrm{HHI}(\mathbf{y}^{(k)})$.
Phase 3: Random restarts and selection
The greedy procedure just described is deterministic conditional on the rule order. To mitigate sensitivity to that order and to escape potential local optima (since different groupings of menus yield different concentrations), we repeat Phase 2 for $K$ different permutations of the rules:
  • iteration $k=1$ uses the popularity–based order (highest–coverage rules first);
  • iterations $k=2,\dots,K$ use independent random permutations of $\mathcal{F}$.
We obtain $K$ candidate assignments $\mathbf{y}^{(1)},\dots,\mathbf{y}^{(K)}$ and define
$\widehat{\operatorname{MRCI}}(D;\mathcal{F})\ :=\ \max_{1\leq k\leq K}\ \mathrm{HHI}(\mathbf{y}^{(k)}).$
5.2 Accuracy and Scalability
The heuristic replaces combinatorial optimization with a sequence of fast linear passes. To validate this approach, we benchmarked the heuristic against the MIQP solver on a set of 3,430 subsampled optimization tasks (see Appendix D for full details).
The results confirm that the heuristic provides a high-quality approximation. The mean absolute difference between the exact and heuristic MRCI is less than $0.001$, and over $96\%$ of estimates fall within a $1\%$ tolerance of the true global optimum. In terms of scalability, the heuristic scales approximately linearly in $T\times\mid\mathcal{F}\mid$ and give typical times, whereas the exact solver’s runtime increases fast with the sample size, reaching already 200 seconds when we run the optimization algorithm in datasets with $140$ observations. Given these results, we treat $\widehat{\operatorname{MRCI}}(D;\mathcal{F})$ as an accurate and scalable proxy for the maximal concentration consistent with local discriminability.
6 Statistical Test of Excess Concentration
This section develops finite-sample inference for the Maximum Rule Concentration Index. The goal is to distinguish genuinely structured “few-rule” behavior from concentration that can arise mechanically under menu-independent randomness.
Setup and test statistic. We fix a subject and define the left-choice indicators and their realized count,
$d_{t}:=\mathbf{1}\{x_{t}=L_{t}^{1}\}\in\{0,1\},\qquad Z:=\sum_{t\in\mathcal{T}}d_{t},\qquad\alpha:=Z/T.$
Given the rule library $\mathcal{F}$ and the strict discriminability sets $\{\mathcal{F}_{t}^{\mathrm{strict}}\}_{t\in\mathcal{T}}$ defined in Section 2.2, the MRCI is the maximal Herfindahl-Hirschman concentration over admissible assignments:
$\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})\;=\;\max_{\mathbf{y}\in\mathcal{C}(D;\geq_{\mathrm{FSD}},\mathcal{F})}\ \sum_{f\in\mathcal{F}}s_{f}(\mathbf{y})^{2},\qquad s_{f}(\mathbf{y}):=\frac{1}{T}\sum_{t\in\mathcal{T}}y_{t,f}.$
Throughout, we keep the subject’s menu sequence $\mathcal{A}=\{A_{t}\}_{t\in\mathcal{T}}$ fixed.
6.1 Random Rule Models and two assumptions
The interpretation of the test is guided by a Random Rule Model (RRM), which makes the source of concentration explicit.
Definition 6 (Random Rule Model).
Fix a library $\mathcal{F}$ that contains the two one-sided attention rules A1 and A2. A Random Rule Model (RRM) consists of:
  • A probability vector $\pi=(\pi_{f})_{f\in\mathcal{F}}\in\Delta^{|\mathcal{F}|-1}$
  • An i.i.d. latent process $(H_{t})_{t\in\mathcal{T}}$ with $\Pr(H_{t}=f)=\pi_{f}$. Given a menu $A_{t}=\{L_{t}^{1},L_{t}^{2}\}$ and a realized latent rule $H_{t}=f$, the agent chooses the option prescribed by $f$ whenever $f$ yields a strict dominance recommendation at $A_{t}$; when $f$ does not strictly discriminate at $A_{t}$, choice is generated by a fixed fallback that does not depend on the realized menu index. [^9] Any fixed fallback is admissible for the interpretation below. In particular, one may take the fallback to be a randomization between A1 and A2 with fixed weights.
In a Random Rule Model, a natural scalar measure of latent rule parsimony is the coincidence probability
$\tau(\pi)\;:=\;\sum_{f\in\mathcal{F}}\pi_{f}^{2},$
the probability that two independent latent draws from $\pi$ coincide. Higher $\tau(\pi)$ corresponds to a more concentrated (few-rule) latent behavior. We now formalize the two assumptions underlying our permutation test.
Assumption $H_{0}$. Conditional on the realized count $Z=\sum_{t\in\mathcal{T}}d_{t}$, the binary vector $(d_{t})_{t\in\mathcal{T}}$ is uniform over all $\{0,1\}$-vectors with exactly $Z$ ones, and is independent of the menu sequence $\{A_{t}\}_{t\in\mathcal{T}}$.
$H_{0}$ is an exchangeability null. Once we condition on the subject’s marginal propensity to pick the left option (summarized by $Z$), the locations of the left choices across menus are as-if random.
Assumption $H_{1}$. Choices are generated by an i.i.d. RRM (Definition 6) with latent propensities $\pi$ that are independent of the menu sequence, and $\tau(\pi)$ is larger than the mechanical attention benchmark induced by $\alpha$, $\tau(\pi)\;>\;\tau_{0}(\alpha)\;:=\;\alpha^{2}+(1-\alpha)^{2}.$.
The benchmark $\tau_{0}(\alpha)$ is the concentration obtained by mixing only the two attention rules A1 and A2 with shares $(\alpha,1-\alpha)$; it is the minimal concentration that is always attainable in our framework (Proposition 2(i)). Assumption $H_{1}$ thus captures the notion that the agent’s rule usage is more concentrated than what can be attributed to an unbalanced left/right marginal alone.
6.2 Permutation distribution and $p$-value
Let $\Pi$ denote the set of all permutations of the index set $\mathcal{T}$. For $\sigma\in\Pi$, define a permuted dataset by keeping menus fixed and reassigning the choice indicators:
$D_{\sigma}\;:=\;\{(\tilde{x}_{t},A_{t})\}_{t\in\mathcal{T}},\qquad\tilde{x}_{t}=\begin{cases}L_{t}^{1},&\text{if }d_{\sigma(t)}=1,\\ L_{t}^{2},&\text{if }d_{\sigma(t)}=0.\end{cases}$
For each $D_{\sigma}$, we recompute local strict discriminability relative to the permuted choices and therefore recompute $\mathcal{C}(D_{\sigma};\geq_{\mathrm{FSD}},\mathcal{F})$ as in Definition 2. This yields the permuted statistic $\operatorname{MRCI}(D_{\sigma};\geq_{\mathrm{FSD}},\mathcal{F})$.
Definition 7.
The (right-tail) permutation $p$-value is
$p_{\mathrm{perm}}(D)\;:=\;\frac{1}{|\Pi|}\sum_{\sigma\in\Pi}\mathbf{1}\!\left\{\operatorname{MRCI}(D_{\sigma};\geq_{\mathrm{FSD}},\mathcal{F})\ \geq\ \operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})\right\}.$
In practice, $|\Pi|=T!$ is enormous. We therefore approximate $p_{\mathrm{perm}}(D)$ by Monte Carlo. We draw $\sigma_{1},\dots,\sigma_{B}$ i.i.d. uniformly from $\Pi$ and compute
$\operatorname{MRCI}^{(b)}:=\operatorname{MRCI}(D_{\sigma_{b}};\geq_{\mathrm{FSD}},\mathcal{F}),\qquad b=1,\dots,B.$
We report the randomized $p$-value with the standard finite-sample correction,
$\widehat{p}_{\mathrm{perm}}(D)\;:=\;\frac{1}{B}\sum_{b=1}^{B}\mathbf{1}\{\operatorname{MRCI}^{(b)}\geq\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})\},$
and reject Assumption $H_{0}$ at level $\eta$ when $\widehat{p}_{\mathrm{perm}}(D)\leq\eta$.
6.3 Statistical Properties
Proposition 4.
Under Assumption $H_{0}$, conditional on $Z$, the permutation test has size $\eta$:
$\Pr\!\big(p_{\mathrm{perm}}(D)\leq\eta\ \big|\ Z\big)\ \leq\ \eta\qquad\text{for all }\eta\in[0,1].$
Consequently, the Monte Carlo version based on $\widehat{p}_{\mathrm{perm}}(D)$ is also finite-sample valid.
Proposition 4 is purely a randomization-test statement: under the exchangeable null, the observed dataset is (conditionally on $Z$) distributed uniformly over its permutation orbit, so the rank of the observed statistic within the orbit is uniform up to ties.
Power. We now state a large-$T$ power property of the permutation test under the excess-concentration alternative.
Assumption 2 (Stabilizing strict-recommendation frequencies).
There exists a sequence of menu collections $\mathcal{A}_{T}=\{A_{t,T}\}_{t=1}^{T}$ with $A_{t,T}=\{L_{t,T}^{1},L_{t,T}^{2}\}$ such that, for every rule $f\in\mathcal{F}$, the fractions of menus at which $f$ yields a strict FSD recommendation for $L^{1}$ and for $L^{2}$ converge:
$s^{(T)}_{f,1}\;:=\;\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\!\left\{\pi_{f}(L_{t,T}^{1};L_{t,T}^{2})\;>_{\mathrm{FSD}}\;\pi_{f}(L_{t,T}^{2};L_{t,T}^{1})\right\}\ \longrightarrow\ s_{f,1}\in[0,1],$
$s^{(T)}_{f,2}\;:=\;\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\!\left\{\pi_{f}(L_{t,T}^{2};L_{t,T}^{1})\;>_{\mathrm{FSD}}\;\pi_{f}(L_{t,T}^{1};L_{t,T}^{2})\right\}\ \longrightarrow\ s_{f,2}\in[0,1].$
$1-s^{(T)}_{f,1}-s^{(T)}_{f,2}$ is the fraction of menus at which $f$ yields no strict recommendation.
Assumption 2 holds, for instance, if menus are i.i.d. draws from a fixed distribution, or if the deterministic design repeats a finite menu set with stable frequencies. Next, to link the observed MRCI to the latent concentration $\tau(\pi)$, we impose that the latent rule draws generate (up to a negligible set of tie events) a compatible assignment in the sense of Local Discriminability.
Assumption 3 (Asymptotically active latent rules).
Under Assumption $H_{1}$, the RRM generates data $D_{T}$ such that the realized latent profile $\mathbf{H}_{T}=(H_{t})_{t=1}^{T}$ is admissible with probability tending to one, i.e.,
$\Pr\!\big(\mathbf{H}_{T}\in\mathcal{C}(D_{T};\geq_{\mathrm{FSD}},\mathcal{F})\big)\ \longrightarrow\ 1.$
Proposition 5 (Consistency under excess concentration).
Let $(D_{T})_{T\geq 1}$ be generated under Assumption $H_{1}$, and suppose that Assumptions 2 and 3 hold. Assume moreover that the left-choice share $\alpha_{T}:=Z_{T}/T$ converges in probability to some $\bar{\alpha}\in(0,1)$, where $Z_{T}=\sum_{t=1}^{T}d_{t}$. If the latent concentration satisfies
$\tau(\pi)\;>\;\tau_{0}(\bar{\alpha})\;=\;\bar{\alpha}^{2}+(1-\bar{\alpha})^{2},$
then the exact permutation $p$-value vanishes:
$p_{\mathrm{perm}}(D_{T})\ \xrightarrow[T\to\infty]{\Pr}\ 0.$
Consequently, for any fixed $\eta\in(0,1)$, the permutation test that rejects when $p_{\mathrm{perm}}(D_{T})\leq\eta$ has asymptotic power one:
$\Pr\!\big(p_{\mathrm{perm}}(D_{T})\leq\eta\big)\xrightarrow[T\to\infty]{\ }1.$
The same conclusion holds for the Monte Carlo version $\widehat{p}_{\mathrm{perm}}(D_{T})$ provided $B\to\infty$.
Proposition 5 formalizes the sense in which the test detects excess concentration. Under a i.i.d. RRM with sufficiently concentrated latent propensities, the observed MRCI is asymptotically bounded below by $\tau(\pi)$, because the latent profile itself is an admissible explanation, while the permutation distribution (which destroys any systematic alignment between menus and choices, conditional on $Z_{T}$) concentrates around the mechanical attention benchmark $\tau_{0}(\alpha_{T})$. When $\tau(\pi)>\tau_{0}(\bar{\alpha})$, these two quantities are separated, so the right-tail permutation probability tends to zero.
Overall, the null hypothesis isolates a strong form of menu-independent randomness. After conditioning on the left-choice count $Z$, the assignment of left/right choices to the realized menus is as-if random. The permutation distribution therefore captures the concentration that can arise from accidental alignment between random left/right choices and the menus that happen to be compatible with a given rule. Rejection indicates that the observed $\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})$ is too large to be plausibly generated by this benchmark. In terms of the RRM interpretation, such rejections are consistent with the presence of excess latent rule concentration. The subject’s choices admit a maximally concentrated rule assignment that is unusually high relative to what would be expected under menu-independent random choice.
7 Empirical Application
Data.
We apply our framework to the publicly available CPC18 competition dataset (erev_ert_plonsky_2017; plonsky2017_cpc18). [^10] The data can be downloaded on the following website: https://cpc-18.com/data/. Since our rule library and dominance benchmark are defined over lotteries with known probabilities, we restrict attention to the risk problems and exclude ambiguity tasks. The resulting sample comprises $N=686$ subjects (average age 24.9, s.d. 2.93; 54% female) recruited across multiple locations in Israel. Each subject faces $20$ to $28$ distinct binary lottery choice problems, and each problem is repeated for $25$ trials, yielding $500$ to $700$ decisions per subject. The design includes both no-feedback and feedback regimes, [^11] In the feedback regime, repeated exposure and realized outcomes may induce learning or history-dependence. and the choice problems were design to elicit canonical behavioral anomalies in risky choice, including Allais paradoxes (allais1953; tversky_kahneman1986), certainty-like effects (kahneman_tversky1979), break-even effects (thaler_johnson1990), splitting effects (birnbaum2008), or systematic departures from expected-value reasoning driven by the weighting of rare events (barron_erev2003). For a detailed description of the experimental design, the exact list of problems, and the mapping from problems to targeted phenomena, see erev_ert_plonsky_2017. Summary statistics are provided in Table 5.
Table 5: Summary Statistics
Statistic Value
Demographics
Number of Subjects 686
Female Share 0.54
Age (Mean) 24.9
Age (Std. Dev.) 2.9
Rule Concentration Measures
MRCI (Mean) 0.545
MRCI (Std. Dev.) 0.05
$N_{eff}$ (Mean) 1.85
Share Single Rule ($N_{eff}\approx 1$) 0
Share Non-Random ($p\leq 0.10$) 0.669
Share Non-Random ($p\leq 0.01$) 0.641
  • Notes: The table reports summary statistics for the $N=686$ subjects in the CPC18 dataset. The lower panel reports the Maximum Rule Concentration Index (MRCI), computed subject-by-subject under the rule library $\mathcal{F}$ and the $\geq_{\mathrm{FSD}}$ benchmark. MRCI is estimated with the heuristic algorithm in Section 5 using $R=100$ independent random restarts per subject; the Effective Number of Rules is $N_{\mathrm{eff}}=1/\operatorname{MRCI}$. The “Share Non-Random” rows report the fraction of subjects for whom we reject the exchangeable random-choice null using the finite-sample permutation test in Section 6. We approximate the exact permutation distribution by Monte Carlo with $B=500$ random permutations. Each permuted-statistic $\operatorname{MRCI}^{(b)}$ is computed using the same heuristic (with $R=100$ restarts).
Implementation.
For each subject $i$, we compute the Maximum Rule Concentration Index $\operatorname{MRCI}(D_{i};\,\geq_{\mathrm{FSD}},\mathcal{F})$ using the optimization routine described in Section 5. In practice, we first pre-compute the menu-by-rule admissibility matrix (which records, for each binary problem faced by the subject, which rules in $\mathcal{F}$ render the realized choice a strict improvement under the FSD benchmark applied to perceived lotteries). We then solve the resulting assignment problem with a greedy-refinement heuristic, and run the algorithm with $100$ independent random restarts. We report two rule-level diagnostics. The concentration gain $\phi(f)$ is computed from “full minus one” comparisons (dropping $f$ from $\mathcal{F}$ while keeping the two one-sided-attention defaults. The stability score $\kappa(f)$ is computed by iteratively attempting to delete rules in random orders and recording how often each rule remains necessary to attain the subject’s maximal concentration; we use $500$ random orderings of the library.
For inference, we implement the finite-sample permutation test described in Section 6. For each subject, we generate the conditional randomization distribution of $\operatorname{MRCI}$ by permuting the realized left/right choices across the subject’s fixed set of menus, and we compute $p$-values using $500$ permutations with 100 independent restarts per permutation. All computations were performed in R, with parallel execution on a 64-core computing cluster; the full empirical analysis completes in under 4 hours.
Results.
We first report the average level of rule concentration. The average MRCI across subjects is $0.55$ (s.d. $0.05$). This corresponds to an average effective number of rules $N_{eff}\approx 1.85$, suggesting that the average subject behaves as if choosing between fewer than two rules. The density plot associated to $N_{eff}$ is provided in Figure 1.
Figure 1: Density of $N_{eff}$
To formally test whether this concentration is spurious, we examine the distribution of $p$-values from our permutation test (Figure 2). We find that the maximal concentration is unusually high relative to what would be expected under menu-independent random choice. 64% of subjects reject the null hypothesis of random choice at the 1% significance level. This confirms that for a substantial share of the population, the observed consistency with the rule library $\mathcal{F}$ cannot be attributed to chance.
Figure 2: Distribution of the p-values from the Permutation Test
Rule Importance.
Which rules drive this concentration? Figure 3 summarizes two complementary notions of importance. The left panel reports the average concentration gain $\phi(f)$, which measures how much maximal concentration falls when rule $f$ is removed from the library. One rule clearly dominates by this metric: MAP (modal payoff focus). On average, MAP accounts for about $17\%$ of attainable concentration, meaning that deleting MAP reduces the subject-level MRCI by roughly $17\%$ of its baseline value. The next contributors are SAL (salience-based focusing) at about $7\%$, followed by MMx, MMn, and REG in the $4$-$5\%$ range. The remaining rules have negligible marginal contributions in terms of maximal concentration.
The right panel reports the stability score $\kappa(f)$, which captures a different notion: how often a rule remains necessary to reach the maximal concentration level along random deletion paths of the library. By this metric, MAP is again most central: it remains essential in roughly $40\%$ of deletion orderings, indicating that many high-concentration explanations rely on MAP -assignments that cannot be fully replicated by reallocating menus to other rules. At the same time, $\kappa(f)$ reveals substantial substitution. Both REG and SAL have sizeable stability (about $25\%$ and $20\%$, respectively), even though their concentration gains are smaller than MAP ’s. This pattern is consistent with REG and SAL acting as structural substitutes: they may not raise maximal concentration as much on average, but they often become pivotal in sustaining a high-concentration assignment once other rules are removed. The extremum heuristics MMx and MMn also exhibit nontrivial stability (roughly $12\%$ down to a few percent), suggesting that they are sometimes required for particular subsets of menus even if they are rarely the primary driver of aggregate concentration.
Finally, three rules are largely inactive by these importance measures, and the interpretation differs across them. For PW (probability weighting) and Id (identity), low $\phi$ and low $\kappa$ do not necessarily mean that these transformations are conceptually irrelevant. They might reflect the fact that these rules, as specified here, primarily re-encode lotteries without reliably generating strict dominance comparisons that would uniquely pin down choices menu-by-menu. In particular, PW distorts probabilities but this distortion alone might not suffice to turn the realized choice into a strict FSD improvement, so it may rarely expand the set of locally admissible assignments. By contrast, DIS (disappointment) appears genuinely non-operative in this dataset under our operationalization: it almost never contributes to maximal concentration and is rarely required to sustain it. We interpret this cautiously as evidence that, within our library and dominance-based discipline, disappointment-based comparisons are not a leading channel for explaining the CPC18 choice patterns.
Figure 3: Rule Importance Metrics. Error bars indicate 95% confidence intervals.
Finally, we investigate heterogeneity across gender. A simple OLS regression of MRCI on gender (controlling for age) reveals a statistically significant difference: women exhibit an MRCI that is approximately 0.01 points lower than men (see Table A.1 in the Appendix). In terms of the effective number of rules, this implies that men are slightly more concentrated in their rule usage, while women use a slightly more diverse set of rules (higher $N_{eff}$). We decompose this difference in Figure 4. The concentration gains for the dominant rules (MAP and SAL) are slightly higher for men, driving their higher overall MRCI. Conversely, women tend to have higher stability scores across a broader range of secondary rules. This supports the interpretation that male subjects in this sample seem to lock onto a slightly narrower set of rules, whereas female subjects display behavior that is better described by a broader weighted portfolio of rules.
Figure 4: Gender Heterogeneity in Rule Usage. The figure compares the Concentration Gain (Panel A) and Stability Score (Panel B) for men (dark blue) and women (orange).
8 Conclusion
This paper develops a revealed-rule approach to risky choice designed for environments in which perception is plausibly menu-dependent. Rather than taking global rationalizability as the primary source of empirical discipline, we assume that at each menu, the rule used must make the realized choice a strict improvement according to a dominance benchmark applied to the perceived lotteries generated by that rule. Under first-order stochastic dominance, this local discriminability requirement has a strong implication: any rule profile that satisfies it induces a perceived dataset that is automatically cycle-free in revealed-preference terms. Empirical content therefore comes from parsimony across menus; how concentrated a rule-based explanation can be.
We quantify this parsimony with the Maximum Rule Concentration Index (MRCI), defined as the maximal Herfindahl-Hirschman concentration of rule shares over all locally admissible menu-by-menu assignments. The MRCI provides a single interpretable summary of behavioral complexity. It is close to one when a small number of mechanisms can repeatedly justify choices across many menus, and it is lower when any admissible explanation must spread weight across a larger portfolio of rules. Because concentration alone does not reveal which mechanisms are responsible, we complement the index with two diagnostics. Concentration gains capture the marginal contribution of each rule to maximal concentration, while stability scores capture how robustly a rule remains necessary to sustain the optimum when the library is pruned. Together, these objects distinguish rules that genuinely unify behavior across many problems from rules that matter mainly as substitutes at the margin.
Methodologically, the framework is built to be usable on standard datasets. Computing the MRCI is an NP-hard assignment problem with a convex objective. We provide a MIQP formulation, but also a scalable heuristic that exploits the structure of concentration maximization. Optimal assignments tend to load heavily on high-coverage rules, and we benchmark it against the exact solution on subsampled instances. For inference, we propose a finite-sample permutation test in the spirit of revealed-preference power analyses. The test is tailored to our setting via a Random Rule Model interpretation: because the library includes one-sided attention defaults, any binary choice sequence is feasible, and under menu-independent randomness the conditional distribution of the MRCI is obtained by permuting the realized left/right choices across fixed menus.
Applied to CPC18, the approach yields a compact characterization of within-subject regularities in a rich battery of repeated lottery choices. On average, subjects behave as if relying on fewer than two effective rules, and more than half reject a menu-independent random benchmark at the one-percent level. The diagnostics sharpen the interpretation: modal-payoff focusing and salience-based focusing account for most of the attainable concentration gains, while regret and extremum heuristics contribute meaningfully but less. At the same time, stability scores indicate substantial substitution among the active rules, consistent with the idea that multiple reduced-form mechanisms can locally justify overlapping subsets of menus even when a small set of “unifiers” largely drives overall concentration. These results illustrate how a large menu-dependent behavioral literature can be brought onto a common footing: not by choosing a single structural model ex ante, but by measuring the minimal rule complexity needed to rationalize observed behavior and identifying which broad mechanisms carry that complexity.
The framework is intentionally modular, and has several limitations. First, our clean implication from local admissibility to strong cyclical consistency relies on first-order stochastic dominance; extending the approach to alternative pre-orders may require integrating local discriminability with additional global constraints. Second, we take the rule library as a primitive. A natural next step is to treat the library itself as an empirical object: to compare competing libraries, to penalize library expansion in a principled way, and to study how identification changes as the menu of candidate mechanisms grows. Third, our analysis is nonstructural about the determinants of rule use. Embedding concentration and rule-importance measures in models with cognitive costs, attention constraints, learning from feedback, incentives, or endogenous consideration would clarify when observed parsimony reflects genuine simplicity versus rational adaptation to the choice architecture. More broadly, the approach suggests an agenda in which decision rules are measured and compared across populations and decision environments, providing a bridge between behavioral theory and revealed preference.
Online Appendix
Appendix A Local Discriminability and Cyclical ConsistencyA.1
To characterize cyclical consistency conditions in our choice environment, we define revealed preference relations relative to the dominance preorder. The following definition follows nishimura2017.
Definition 8.
[…]
    $x=x_{t}\quad\text{and}\quad y\in A_{t}^{\downarrow},$
    where
    $A_{t}^{\downarrow}:=\{\,y\in X:\ \exists\,z\in A_{t}\text{ with }z\trianglerighteq y\,\}$
[…]
  • We say that $x$ is directly revealed strictly $\triangleright$–preferred to $y$, and write $x\mathrel{P}y$, if there exists $t\in\mathcal{T}$ such that
    $x=x_{t}\quad\text{and}\quad y\in A_{t}^{\downarrow\downarrow},$
    where
    $A_{t}^{\downarrow\downarrow}:=\{\,y\in X:\ \exists\,z\in A_{t}\text{ with }z\triangleright y\,\}$
[…]
  • $R^{\ast}$ denotes the transitive closure of $R$.
  • $P^{\ast}$ denotes the strict transitive closure induced by $(R,P)$: we write $x\mathrel{P^{\ast}}y$ if there exists a finite sequence $x=x^{(1)},x^{(2)},\dots,x^{(M)}=y$ such that each step is either $x^{(m)}Rx^{(m+1)}$ or $x^{(m)}Px^{(m+1)}$, and at least one step is strict (of type $P$).
Thus $xR^{\ast}y$ means that there is a finite sequence $x=x^{(1)},x^{(2)},\dots,x^{(M)}=y$ such that each step $x^{(m)}Rx^{(m+1)}$ is justified by some observation in $D$. Moreover, $xP^{\ast}y$ means that there is such a sequence with at least one strict step (of type $P$).
Weak Cyclical Consistency. In the standard consumer–demand setting, the Generalized Axiom of Revealed Preference (GARP) rules out revealed preference cycles that contain at least one strict step. The analogue in our framework is defined below.
[…]
$x\mathrel{R^{\ast}}y\quad\text{and}\quad y\mathrel{P}x.$
[…]
$x\mathrel{R^{\ast}}y\quad\text{and}\quad y\mathrel{R^{\ast}}x.$
Rationality benchmark.
[…]
Proposition A.1.
Suppose $\trianglerighteq=\geq_{\mathrm{FSD}}$, and let $\mathbf{f}\in\mathcal{F}^{\mathcal{T}}$ satisfy Assumption 1. Then the perceived dataset $\widetilde{D}(\mathbf{f})=\{(\widetilde{x}_{t}(\mathbf{f}),\widetilde{A}_{t}(\mathbf{f}))\}_{t\in\mathcal{T}}$ satisfies strong cyclical $\geq_{\mathrm{FSD}}$–consistency (Definition 10).
The proof is provided below. The core intuition is simple and relies on an order-theoretic feature of First-Order Stochastic Dominance. Local Discriminability forces each applied rule to turn the observed choice into a clear improvement in the perceived menu: at every menu, the perceived chosen lottery is strictly better than its perceived rival according to FSD. Because we define revealed preference using the FSD–downward closure of each perceived menu, this local strictness has a strong spillover: whenever an alternative is deemed “available” for revealed preference comparisons (i.e., it lies weakly below one of the two perceived menu elements), it is necessarily weakly below the perceived chosen element as well. In other words, in the perceived dataset, each menu points “downwards” in an FSD sense from the perceived chosen lottery. As a result, any chain of revealed preference steps in $\widetilde{D}(\mathbf{f})$ can only move in one direction with respect to FSD. A nontrivial cycle would require moving down and eventually coming back up to a distinct lottery, which is impossible because FSD does not allow strict improvements to be undone: if two lotteries weakly dominate each other under FSD, they must coincide. This is the sense in which strong cyclical $\geq_{FSD}$-consistency in the perceived dataset is not an additional restriction but a consequence of local rule-based justification under FSD.
Proof.
Fix a rule profile $\mathbf{f}\in\mathcal{F}^{T}$ satisfying Local Discriminability, i.e., for every observation $t$ with menu $\widetilde{A}_{t}(\mathbf{f})=\{\widetilde{L}^{1}_{t,f_{t}},\widetilde{L}^{2}_{t,f_{t}}\}$ and perceived choice $\widetilde{x}_{t}(\mathbf{f})=\widetilde{L}^{i_{t}}_{t,f_{t}}$, we have
$\widetilde{L}^{i_{t}}_{t,f_{t}}\;>_{\mathrm{FSD}}\;\widetilde{L}^{-i_{t}}_{t,f_{t}}.$
For brevity, in what follows we suppress $\mathbf{f}$ from the notation and write $\widetilde{x}_{t}$ for the perceived chosen lottery at $t$ and $\widetilde{y}_{t}$ for its perceived alternative, so that
$\widetilde{x}_{t}\;>_{\mathrm{FSD}}\;\widetilde{y}_{t}\qquad\text{for all }t\in\mathcal{T}.$
We work with the revealed preference relations defined on the perceived dataset $\widetilde{D}(\mathbf{f})$.
  • For a menu $\widetilde{A}_{t}$, its FSD–downward closure is
    $\widetilde{A}_{t}^{\downarrow}\;:=\;\{\,z\in X:\exists\,w\in\widetilde{A}_{t}\text{ with }w\geq_{\mathrm{FSD}}z\,\}.$
  • The direct revealed preference relation $\widetilde{R}$ on $X$ is defined by
    $x\,\widetilde{R}\,y\quad\Longleftrightarrow\quad\exists\,t\in\mathcal{T}\text{ such that }x=\widetilde{x}_{t}\text{ and }y\in\widetilde{A}_{t}^{\downarrow}.$
  • The transitive closure $\widetilde{R}^{\ast}$ is defined in the usual way: $x\,\widetilde{R}^{\ast}\,y$ if and only if there exists a finite sequence $x_{0},\dots,x_{n}$ with $x_{0}=x$, $x_{n}=y$ and $x_{k}\,\widetilde{R}\,x_{k+1}$ for all $k=0,\dots,n-1$.
We must show that $\widetilde{R}^{\ast}$ has no non-trivial cycle, i.e., there do not exist $u,v\in X$ with $u\neq v$ such that
$u\,\widetilde{R}^{\ast}\,v\quad\text{and}\quad v\,\widetilde{R}^{\ast}\,u.$
Step 1: Direct revealed preference is FSD–monotone.
We first prove that whenever $x\,\widetilde{R}\,y$ holds, we have
$x\;\geq_{\mathrm{FSD}}\;y.$
Take any pair $(x,y)$ with $x\,\widetilde{R}\,y$. By definition of $\widetilde{R}$, there exists some $t\in\mathcal{T}$ such that
$x=\widetilde{x}_{t}\quad\text{and}\quad y\in\widetilde{A}_{t}^{\downarrow}.$
By definition of $\widetilde{A}_{t}^{\downarrow}$, there exists $w\in\widetilde{A}_{t}$ with
$w\;\geq_{\mathrm{FSD}}\;y.$
The menu $\widetilde{A}_{t}$ contains exactly the two lotteries $\widetilde{x}_{t}$ and $\widetilde{y}_{t}$, so $w$ is either $\widetilde{x}_{t}$ or $\widetilde{y}_{t}$.
Case 1: $w=\widetilde{x}_{t}$. Then $\widetilde{x}_{t}\geq_{\mathrm{FSD}}y$ holds directly, so $x=\widetilde{x}_{t}\geq_{\mathrm{FSD}}y$.
Case 2: $w=\widetilde{y}_{t}$. By Local Discriminability, we have
$\widetilde{x}_{t}\;>_{\mathrm{FSD}}\;\widetilde{y}_{t},$
which implies $\widetilde{x}_{t}\geq_{\mathrm{FSD}}\widetilde{y}_{t}$. Combining this with $w=\widetilde{y}_{t}\geq_{\mathrm{FSD}}y$ and transitivity of $\geq_{\mathrm{FSD}}$, we obtain
$\widetilde{x}_{t}\;\geq_{\mathrm{FSD}}\;y.$
Thus again $x=\widetilde{x}_{t}\geq_{\mathrm{FSD}}y$. In both cases we conclude that $x\geq_{\mathrm{FSD}}y$ whenever $x\,\widetilde{R}\,y$.
Step 2: Any revealed–preference path is FSD–monotone.
Let $x,y\in X$ with $x\,\widetilde{R}^{\ast}\,y$. By definition of $\widetilde{R}^{\ast}$, there exist $n\geq 1$ and a sequence $x_{0},\dots,x_{n}$ in $X$ such that
$x_{0}=x,\quad x_{n}=y,\quad\text{and}\quad x_{k}\,\widetilde{R}\,x_{k+1}\ \text{ for all }k=0,\dots,n-1.$
From Step 1, each direct edge $x_{k}\,\widetilde{R}\,x_{k+1}$ implies
$x_{k}\;\geq_{\mathrm{FSD}}\;x_{k+1}.$
Using transitivity of $\geq_{\mathrm{FSD}}$ repeatedly along the path $(x_{k})_{k=0}^{n}$, we obtain
$x=x_{0}\;\geq_{\mathrm{FSD}}\;x_{1}\;\geq_{\mathrm{FSD}}\;\dots\;\geq_{\mathrm{FSD}}\;x_{n}=y,$
hence
$x\;\geq_{\mathrm{FSD}}\;y.$
Thus we have shown:
$x\,\widetilde{R}^{\ast}\,y\quad\Longrightarrow\quad x\geq_{\mathrm{FSD}}y.$
Step 3: Absence of strong cycles.
Suppose, for contradiction, that there exist two distinct lotteries $u,v\in X$ with
$u\,\widetilde{R}^{\ast}\,v\quad\text{and}\quad v\,\widetilde{R}^{\ast}\,u,\quad\text{and }u\neq v.$
Applying the implication from Step 2 twice, we obtain
$u\;\geq_{\mathrm{FSD}}\;v\quad\text{and}\quad v\;\geq_{\mathrm{FSD}}\;u.$
The FSD relation $\geq_{\mathrm{FSD}}$ is a partial order on $X$: it is reflexive, transitive, and antisymmetric. In particular, antisymmetry states that
$\big(u\geq_{\mathrm{FSD}}v\text{ and }v\geq_{\mathrm{FSD}}u\big)\quad\Longrightarrow\quad u=v.$
Hence the two inequalities above imply $u=v$, contradicting the assumption $u\neq v$. Therefore no such pair $(u,v)$ can exist, and the perceived dataset $\widetilde{D}(\mathbf{f})$ satisfies strong cyclical $\geq_{\mathrm{FSD}}$–consistency. ∎
Appendix B Proofs
B.1 Proof of Proposition 2
Let $T$ be the number of menus, $s_{f}(\mathbf{y}):=\frac{1}{T}\sum_{t=1}^{T}y_{t,f}$ and $\mathcal{F}^{\star}(\mathbf{y}):=\{f\in\mathcal{F}:\ s_{f}(\mathbf{y})>0\}$.
(i) Bounds. Fix any feasible profile $\mathbf{y}\in\mathcal{C}(D)$ and let $s_{f}(\mathbf{y})\geq 0$ denote the induced rule shares with $\sum_{f\in\mathcal{F}}s_{f}(\mathbf{y})=1$. By definition,
$\mathrm{HHI}(\mathbf{y})=\sum_{f\in\mathcal{F}}s_{f}(\mathbf{y})^{2}.$
Upper bound: since $s_{f}(\mathbf{y})\in[0,1]$ and the shares sum to one, we have $\sum_{f}s_{f}^{2}\leq 1$, with equality only when one share equals $1$ and all others are 0. Hence $\mathrm{HHI}(\mathbf{y})\leq 1$ for every feasible $\mathbf{y}$, and therefore $\mathrm{MRCI}=\max_{\mathbf{y}\in\mathcal{C}(D)}\mathrm{HHI}(\mathbf{y})\leq 1$. The proof for the lower bound is provided in the text.
(ii) $\mathrm{MRCI}=1$ iff a one–rule rationalization exists. If $\mathrm{MRCI}=1$, let $\mathbf{y}^{\star}$ be an optimizer. Since $\sum_{f}s_{f}(\mathbf{y}^{\star})=1$ and $\sum_{f}s_{f}(\mathbf{y}^{\star})^{2}=1$, it must be that $s_{f^{\dagger}}(\mathbf{y}^{\star})=1$ for a single $f^{\dagger}$ and $s_{f}(\mathbf{y}^{\star})=0$ otherwise.
(iii) Replication invariance. Form $D^{(k)}$ by stacking $k$ copies of $D$, so $T^{(k)}=kT$. Replicate any feasible assignment $\mathbf{y}$ blockwise to obtain $\mathbf{y}^{(k)}$. The set $\mathcal{F}_{t}^{\rm strict}$ repeat in each block, and local discriminability is also repeated across blocks. Thus $\mathbf{y}^{(k)}\in\mathcal{C}(D^{(k)})$. Let $n_{f}=\sum_{t}y_{t,f}$ and $n_{f}^{(k)}=\sum_{t^{\prime}}y^{(k)}_{t^{\prime},f}=k\,n_{f}$. Then
$\mathrm{HHI}(\mathbf{y}^{(k)})=\sum_{f}\Big(\frac{n^{(k)}_{f}}{T^{(k)}}\Big)^{2}=\sum_{f}\Big(\frac{k\,n_{f}}{k\,T}\Big)^{2}=\sum_{f}\Big(\frac{n_{f}}{T}\Big)^{2}=\mathrm{HHI}(\mathbf{y}).$
Therefore the feasible objective set is preserved, implying $\mathrm{MRCI}(D^{(k)})=\mathrm{MRCI}(D)$.
(4) Forced choice under unique local discriminability. Fix $t$ with $\mathcal{F}_{t}^{\rm strict}=\{f^{\ast}\}$. By Assumption 1, feasibility at menu $t$ requires selecting a strictly discriminating rule whenever one exists; thus any $\mathbf{y}\in\mathcal{C}(D)$ must satisfy $y_{t,f^{\ast}}=1$ and $y_{t,f}=0$ for all $f\neq f^{\ast}$. Hence every optimizer necessarily uses $f^{\ast}$ at $t$.
B.2 Proof of Proposition 3
Monotonicity. Let $y^{\star}\in\mathcal{C}(D;\mathcal{F})$ attain $\mathrm{MRCI}(D;\geq_{FSD},\mathcal{F})$. Define $y^{\prime}\in\mathcal{C}(D;\mathcal{F}\cup\{f\})$ by $y^{\prime}_{t,r}=y^{\star}_{t,r}$ for $r\in\mathcal{F}$ and $y^{\prime}_{t,f}=0$ for all $t$. Then $\mathrm{HHI}(y^{\prime})=\mathrm{HHI}(y^{\star})$, hence the maximum under the larger feasible set is weakly higher.
Necessary condition for strictness. Fix any $\mathcal{F}$–optimizer $y^{\star}$ with shares $s^{\star}=(s_{r}^{\star})_{r\in\mathcal{F}}$ and set $s_{f}^{\star}=0$. Consider any $\mathcal{F}\cup\{f\}$–feasible reassignment that moves fractions $\delta_{r}\in[0,s_{r}^{\star}]$ from rules $r\in\mathcal{F}$ to $f$, producing new shares
$s_{f}^{\prime}=\sum_{r}\delta_{r},\qquad s_{r}^{\prime}=s_{r}^{\star}-\delta_{r}\quad(r\in\mathcal{F}).$
Feasibility requires menu–level admissibility for each reassigned observation (Local Discriminability). The Herfindahl change relative to $s_{f}^{\star}=0$ is
$\Delta\mathrm{HHI}=\Big(\sum_{r}\delta_{r}\Big)^{2}+\sum_{r}(s_{r}^{\star}-\delta_{r})^{2}-\sum_{r}(s_{r}^{\star})^{2}=(\sum_{r}\delta_{r})^{2}\ -\ 2\sum_{r}s_{r}^{\star}\delta_{r}\ +\ \sum_{r}\delta_{r}^{2}.$ (A.1)
If mass is taken from a single donor $r$ ($\delta_{r}>0$, $\delta_{q\neq r}=0$), then
$\Delta\mathrm{HHI}=2\,\delta_{r}(\delta_{r}-s_{r}^{\star})\ \leq\ 0,$
with equality only for $\delta_{r}\in\{0,s_{r}^{\star}\}$ (no move or full relabeling). Hence any one–donor reassignment cannot strictly raise concentration at $s_{f}^{\star}=0$.
Therefore, a necessary condition for strict improvement over $\mathrm{HHI}(y^{\star})$ is: there exist two rules $r\neq s$ with $s_{r}^{\star},s_{s}^{\star}>0$ and a feasible reassignment with $\delta_{r}>0$ and $\delta_{s}>0$. Feasibility, in turn, requires that $f$ be locally admissible on at least one menu currently assigned to $r$ and on at least one menu currently assigned to $s$. If $f$ is never locally admissible then all feasible moves satisfy $\Delta\mathrm{HHI}\leq 0$, and the maximum under $\mathcal{F}\cup\{f\}$ cannot strictly exceed that under $\mathcal{F}$. Since the argument applies to every $\mathcal{F}$–optimizer $y^{\star}$, the stated condition is necessary for strict inequality of the maxima.
B.3 Proof of Proposition 4
Fix the realized menu sequence $\{A_{t}\}_{t\in\mathcal{T}}$ and condition on the realized count $Z=\sum_{t\in\mathcal{T}}d_{t}$. Under Assumption $H_{0}$, the vector $d:=(d_{t})_{t\in\mathcal{T}}$ is uniform over the finite set
$\mathcal{D}_{Z}\;:=\;\Big\{(b_{t})_{t\in\mathcal{T}}\in\{0,1\}^{T}:\ \sum_{t\in\mathcal{T}}b_{t}=Z\Big\},$
and is independent of the menus. Let $\sigma$ be drawn uniformly from the permutation group $\Pi$, independently of the data. Then $d\circ\sigma:=(d_{\sigma(t)})_{t\in\mathcal{T}}$ is also uniform on $\mathcal{D}_{Z}$, and in particular
$(d_{t})_{t\in\mathcal{T}}\ \overset{d}{=}\ (d_{\sigma(t)})_{t\in\mathcal{T}}\qquad\text{conditional on }Z.$
Because $D$ and $D_{\sigma}$ are constructed from the same fixed menus by the deterministic map $b\mapsto\{(x_{t}(b),A_{t})\}_{t\in\mathcal{T}}$ with $x_{t}(b)=L_{t}^{1}$ if $b_{t}=1$ and $x_{t}(b)=L_{t}^{2}$ otherwise, it follows that
$D\ \overset{d}{=}\ D_{\sigma}\qquad\text{conditional on }Z.$
Let $T(\cdot):=\operatorname{MRCI}(\cdot;\geq_{\mathrm{FSD}},\mathcal{F})$ be the statistic. Conditional on $(Z,\{A_{t}\})$, the multiset $\{T(D_{\sigma})\}_{\sigma\in\Pi}$ is well-defined, and $T(D)$ is distributed as a uniform draw from this multiset (up to ties). The permutation $p$-value
$p_{\mathrm{perm}}(D)=\frac{1}{|\Pi|}\sum_{\sigma\in\Pi}\mathbf{1}\{T(D_{\sigma})\geq T(D)\}$
is therefore the usual randomization-test $p$-value for the right tail. Standard randomization-test arguments imply that the conditional distribution of $p_{\mathrm{perm}}(D)$ is stochastically larger than $\mathrm{Unif}[0,1]$ (with equality in the absence of ties), hence
$\Pr\!\big(p_{\mathrm{perm}}(D)\leq\eta\ \big|\ Z\big)\ \leq\ \eta\qquad\forall\,\eta\in[0,1].$
For the Monte Carlo version, the same conclusion follows: conditional on $(D,Z)$, $\widehat{p}_{\mathrm{perm}}(D)$ is the randomization $p$-value for the experiment that draws one “observed” statistic and $B$ i.i.d. “placebo” statistics from the same randomization distribution.
B.4 Proof of Proposition 5
Write $T(\cdot):=\operatorname{MRCI}(\cdot;\geq_{\mathrm{FSD}},\mathcal{F})$ and $\tau_{0}(\alpha):=\alpha^{2}+(1-\alpha)^{2}$. Recall $\alpha_{T}:=Z_{T}/T$ where $Z_{T}=\sum_{t=1}^{T}d_{t,T}$.
Roadmap. Step 1 lower bounds the observed statistic by the latent concentration. Step 2 upper bounds the permutation statistic by $\tau_{0}(\alpha_{T})+o_{p}(1)$, and is the main non-obvious part: (i) we define precisely on which menus a rule can be assigned under strict discriminability, including the possibility that a rule makes no strict recommendation; (ii) we translate this into a cap on the empirical share of each rule; (iii) we prove a deterministic lemma showing that such caps imply an upper bound on the maximal concentration; (iv) we control the caps under random permutation using sampling-without-replacement concentration. Step 3 combines separation to show $p_{\mathrm{perm}}(D_{T})\to 0$.
Step 1 (Observed statistic is bounded below by latent concentration). Let $N_{f,T}:=\sum_{t=1}^{T}\mathbf{1}\{H_{t}=f\}$ and $\widehat{\pi}_{f,T}:=N_{f,T}/T$ be the empirical frequencies of latent rule draws. By Assumption 3, with probability tending to one the latent profile $\mathbf{H}_{T}$ is an admissible assignment for $D_{T}$. Therefore, on that event,
$T(D_{T})\ \geq\ \operatorname{HHI}(\mathbf{H}_{T})\ =\ \sum_{f\in\mathcal{F}}\widehat{\pi}_{f,T}^{\,2}.$
Since $(H_{t})$ are i.i.d. with distribution $\pi$, we have $\widehat{\pi}_{f,T}\to\pi_{f}$ almost surely for each $f$, hence $\sum_{f}\widehat{\pi}_{f,T}^{\,2}\to\sum_{f}\pi_{f}^{2}=\tau(\pi)$ almost surely. Thus,
$T(D_{T})\ \geq\ \tau(\pi)+o_{p}(1).$ (A.2)
Step 2 (Permutation statistic is asymptotically bounded by $\tau_{0}(\alpha_{T})$). Fix $T$ and condition throughout Step 2 on the realized menu sequence $\mathcal{A}_{T}=(A_{t,T})_{t=1}^{T}$ and on $Z_{T}$ (equivalently on $\alpha_{T}$). Let $\sigma$ be a uniform random permutation of $\{1,\dots,T\}$ drawn conditional on preserving $Z_{T}$, and let $D_{T,\sigma}$ be the permuted dataset, i.e. with permuted choice indicators $\tilde{d}_{t,T}:=d_{\sigma(t),T}$.
Step 2a (Where can a rule be assigned?). Fix a rule $f\in\mathcal{F}$. Because $f$ need not generate a strict ordering on every menu, partition the periods into three sets (these depend only on $(A_{t,T},f)$):
$\mathcal{T}^{1}_{f,T}:=\Big\{t:\ \pi_{f}(L_{t,T}^{1};L_{t,T}^{2})\ >_{\mathrm{FSD}}\ \pi_{f}(L_{t,T}^{2};L_{t,T}^{1})\Big\},$
$\mathcal{T}^{2}_{f,T}:=\Big\{t:\ \pi_{f}(L_{t,T}^{2};L_{t,T}^{1})\ >_{\mathrm{FSD}}\ \pi_{f}(L_{t,T}^{1};L_{t,T}^{2})\Big\},\qquad\mathcal{T}^{0}_{f,T}:=\{1,\dots,T\}\setminus(\mathcal{T}^{1}_{f,T}\cup\mathcal{T}^{2}_{f,T}).$
On $\mathcal{T}^{0}_{f,T}$ the rule makes no strict recommendation (tie/inaction), hence it cannot be assigned under strict discriminability regardless of the realized choice.
Define indicators $a^{(1)}_{t,T}(f):=\mathbf{1}\{t\in\mathcal{T}^{1}_{f,T}\}$ and $a^{(2)}_{t,T}(f):=\mathbf{1}\{t\in\mathcal{T}^{2}_{f,T}\}$, so $a^{(1)}_{t,T}(f)+a^{(2)}_{t,T}(f)\leq 1$.
Step 2b (Admissible assignments imply share caps). An assignment is a matrix $\mathbf{y}=(y_{t,f})_{t\leq T,\,f\in\mathcal{F}}$ with $y_{t,f}\in\{0,1\}$ and $\sum_{f\in\mathcal{F}}y_{t,f}=1$ for every $t$. It is admissible for $D_{T,\sigma}$ if whenever $y_{t,f}=1$, rule $f$ strictly recommends the realized permuted choice at menu $A_{t,T}$. Since the realized permuted choice is $L_{t,T}^{1}$ iff $\tilde{d}_{t,T}=1$ and is $L_{t,T}^{2}$ iff $\tilde{d}_{t,T}=0$, admissibility implies the pointwise bound.
$y_{t,f}\ \leq\ a^{(1)}_{t,T}(f)\,\tilde{d}_{t,T}\ +\ a^{(2)}_{t,T}(f)\,(1-\tilde{d}_{t,T}),\qquad t=1,\dots,T.$ (A.3)
(If $t\in\mathcal{T}^{0}_{f,T}$, then $a^{(1)}_{t,T}(f)=a^{(2)}_{t,T}(f)=0$ and the bound forces $y_{t,f}=0$.)
Define the empirical share of rule $f$ under assignment $\mathbf{y}$ by
$s_{f}(\mathbf{y}):=\frac{1}{T}\sum_{t=1}^{T}y_{t,f}.$
Summing (A.3) over $t$ and dividing by $T$ yields, for every admissible $\mathbf{y}$ and every $f$,
$s_{f}(\mathbf{y})\ \leq\ \frac{1}{T}\sum_{t=1}^{T}\Big(a^{(1)}_{t,T}(f)\,\tilde{d}_{t,T}+a^{(2)}_{t,T}(f)\,(1-\tilde{d}_{t,T})\Big)\ :=\ M_{f,T}(\sigma).$ (A.4)
Thus $M_{f,T}(\sigma)$ is the maximum feasible share that rule $f$ can ever attain under any admissible assignment, given the permuted choices.
Let $m_{T}(\sigma):=\max_{f\in\mathcal{F}}M_{f,T}(\sigma)$.
Step 2c (Deterministic lemma: share caps bound maximal concentration). We now use a purely algebraic fact: if each coordinate of a share vector is capped by $m$, then the Herfindahl index is maximized by placing mass $m$ on one coordinate and the remaining mass 1m1-m on a second coordinate.
Lemma 1 (Cap lemma, case $m\geq 1/2$).
Let $s=(s_{1},\dots,s_{n})$ satisfy $s_{i}\geq 0$, $\sum_{i=1}^{n}s_{i}=1$, and $s_{i}\leq m$ for all $i$, for some $m\in[1/2,1]$. Then
$\sum_{i=1}^{n}s_{i}^{2}\ \leq\ m^{2}+(1-m)^{2}.$
Proof.
We maximize $\sum_{i}s_{i}^{2}$ over the feasible set. Since the objective is strictly convex, any maximizer occurs at an extreme point of the polytope $\{s:\sum_{i}s_{i}=1,\ 0\leq s_{i}\leq m\}$.
Claim 1: at a maximizer, some coordinate equals $m$. Suppose for contradiction that at a maximizer all coordinates satisfy $s_{i}<m$. Because $\sum_{i}s_{i}=1$, there exist indices $j\neq k$ with $s_{k}>0$. For $\varepsilon>0$ small, define $s^{\prime}_{j}=s_{j}+\varepsilon$, $s^{\prime}_{k}=s_{k}-\varepsilon$, and $s^{\prime}_{i}=s_{i}$ otherwise. This keeps $\sum_{i}s^{\prime}_{i}=1$ and preserves feasibility for small enough $\varepsilon$ (since $s_{j}<m$ and $s_{k}>0$). But
$\sum_{i}(s^{\prime}_{i})^{2}-\sum_{i}s_{i}^{2}=(s_{j}+\varepsilon)^{2}+(s_{k}-\varepsilon)^{2}-(s_{j}^{2}+s_{k}^{2})=2\varepsilon(s_{j}-s_{k})+2\varepsilon^{2}.$
Choosing $j$ with $s_{j}\geq s_{k}$ makes the increment strictly positive, contradicting optimality. Hence at least one coordinate must hit the upper bound: $\max_{i}s_{i}=m$.
Claim 2: given one coordinate equals $m$, the remainder should be concentrated on a single other coordinate. Let $s_{1}=m$ wlog, and let r:=1mr:=1-m be the remaining mass. Because $m\geq 1/2$, we have $r\leq m$, so the vector $(m,r,0,\dots,0)$ is feasible. For any feasible allocation of the remainder among coordinates $2,\dots,n$, Cauchy–Schwarz gives
$\sum_{i=2}^{n}s_{i}^{2}\ \leq\ \Big(\sum_{i=2}^{n}s_{i}\Big)^{2}\ =\ r^{2},$
with equality iff all the remainder $r$ is placed on a single coordinate. Therefore,
$\sum_{i=1}^{n}s_{i}^{2}=m^{2}+\sum_{i=2}^{n}s_{i}^{2}\ \leq\ m^{2}+r^{2}=m^{2}+(1-m)^{2},$
as claimed. ∎
Lemma 2 (A built-in lower bound on the maximal match share).
Fix $T$ and condition on $(\mathcal{A}_{T},Z_{T})$, where $\mathcal{A}_{T}=\{A_{t,T}\}_{t=1}^{T}$ and $Z_{T}:=\sum_{t=1}^{T}d_{t,T}$ with $\alpha_{T}:=Z_{T}/T$. Suppose the rule class $\mathcal{F}$ contains the two attention rules $A_{1}$ and $A_{2}$ such that for every menu $A_{t,T}$ they strictly recommend, respectively, $L^{1}_{t,T}$ and $L^{2}_{t,T}$:
$\pi_{A_{1}}(L^{1}_{t,T};L^{2}_{t,T})\;>_{\mathrm{FSD}}\;\pi_{A_{1}}(L^{2}_{t,T};L^{1}_{t,T})\quad\text{and}\quad\pi_{A_{2}}(L^{2}_{t,T};L^{1}_{t,T})\;>_{\mathrm{FSD}}\;\pi_{A_{2}}(L^{1}_{t,T};L^{2}_{t,T}),\qquad t=1,\dots,T.$
(Equivalently, for all $t$: $a^{(1)}_{t,T}(A_{1})=1$, $a^{(2)}_{t,T}(A_{1})=0$, and $a^{(2)}_{t,T}(A_{2})=1$, $a^{(1)}_{t,T}(A_{2})=0$.) Then, for every permutation $\sigma$,
$m_{T}(\sigma)\ :=\ \max_{f\in\mathcal{F}}M_{f,T}(\sigma)\ \geq\ \max\{\alpha_{T},1-\alpha_{T}\}\ \geq\ \tfrac{1}{2}.$
Proof.
Recall
$M_{f,T}(\sigma)=\frac{1}{T}\sum_{t=1}^{T}\Big(a^{(1)}_{t,T}(f)\,d_{\sigma(t),T}\;+\;a^{(2)}_{t,T}(f)\,(1-d_{\sigma(t),T})\Big).$
Because $A_{1}$ strictly recommends $L^{1}_{t,T}$ at every menu, we have $a^{(1)}_{t,T}(A_{1})=1$ and $a^{(2)}_{t,T}(A_{1})=0$ for all $t$, so
$M_{A_{1},T}(\sigma)=\frac{1}{T}\sum_{t=1}^{T}d_{\sigma(t),T}=\frac{1}{T}\sum_{t=1}^{T}d_{t,T}=\alpha_{T},$
where the second equality uses that $\sigma$ is a permutation. Similarly, since $A_{2}$ strictly recommends $L^{2}_{t,T}$ at every menu, $a^{(2)}_{t,T}(A_{2})=1$ and $a^{(1)}_{t,T}(A_{2})=0$ for all $t$, hence
$M_{A_{2},T}(\sigma)=\frac{1}{T}\sum_{t=1}^{T}(1-d_{\sigma(t),T})=1-\alpha_{T}.$
Therefore,
$m_{T}(\sigma)=\max_{f\in\mathcal{F}}M_{f,T}(\sigma)\ \geq\ \max\{M_{A_{1},T}(\sigma),M_{A_{2},T}(\sigma)\}=\max\{\alpha_{T},1-\alpha_{T}\}\ \geq\ \tfrac{1}{2},$
as claimed. ∎
From Lemma 2, we can apply Lemma 1. Applying Lemma 1 to the share vector $(s_{f}(\mathbf{y}))_{f\in\mathcal{F}}$ under any admissible assignment $\mathbf{y}$, using the cap $s_{f}(\mathbf{y})\leq m_{T}(\sigma)$ implied by (A.4), gives
$\sum_{f\in\mathcal{F}}s_{f}(\mathbf{y})^{2}\ \leq\ m_{T}(\sigma)^{2}+\big(1-m_{T}(\sigma)\big)^{2}.$
Taking the maximum over admissible assignments (the definition of $T(D_{T,\sigma})$) yields the key inequality
$T(D_{T,\sigma})\ \leq\ m_{T}(\sigma)^{2}+\big(1-m_{T}(\sigma)\big)^{2}.$ (A.5)
Step 2d (Control $m_{T}(\sigma)$ under random permutation). Define the menu-frequency quantities
$p^{1}_{f,T}:=\frac{1}{T}\sum_{t=1}^{T}a^{(1)}_{t,T}(f)=\frac{|\mathcal{T}^{1}_{f,T}|}{T},\qquad p^{2}_{f,T}:=\frac{1}{T}\sum_{t=1}^{T}a^{(2)}_{t,T}(f)=\frac{|\mathcal{T}^{2}_{f,T}|}{T}.$
By construction $p^{1}_{f,T},p^{2}_{f,T}\in[0,1]$ and $p^{1}_{f,T}+p^{2}_{f,T}\leq 1$.
Using (A.4) and writing $\tilde{d}_{t,T}=d_{\sigma(t),T}$,
$M_{f,T}(\sigma)=\frac{1}{T}\sum_{t\in\mathcal{T}^{1}_{f,T}}\tilde{d}_{t,T}+\frac{1}{T}\sum_{t\in\mathcal{T}^{2}_{f,T}}(1-\tilde{d}_{t,T}).$
Conditional on $(\mathcal{A}_{T},Z_{T})$, the permuted vector $(\tilde{d}_{t,T})_{t=1}^{T}$ is a simple random sample without replacement with exactly $Z_{T}$ ones. Therefore each sum above is (scaled) hypergeometric [^1] Conditional on $(\mathcal{A}_{T},Z_{T})$, the permuted vector $\tilde{d}_{t,T}=d_{\sigma(t),T}$ is a uniform random rearrangement of a length-$T$ $\{0,1\}$ vector with exactly $Z_{T}$ ones. Hence, for any fixed index set $S\subset\{1,\dots,T\}$ with $|S|=n$, the count $X_{S}:=\sum_{t\in S}\tilde{d}_{t,T}$ equals the number of ones observed when drawing $n$ elements without replacement from a population of size $T$ containing $Z_{T}$ ones; therefore $X_{S}\sim\mathrm{Hypergeom}(T,Z_{T},n)$ and $\mathbb{E}[X_{S}]=n(Z_{T}/T)=n\alpha_{T}$. Moreover, by standard concentration/LLN for sampling without replacement (equivalently, for hypergeometric draws), $X_{S}/n\to_{p}Z_{T}/T$ as $n\to\infty$, so $\frac{1}{T}X_{S}=\alpha_{T}\,(n/T)+o_{p}(1)$. , and a sampling-without-replacement law of large numbers implies, for each fixed $f$,
$\frac{1}{T}\sum_{t\in\mathcal{T}^{1}_{f,T}}\tilde{d}_{t,T}=\alpha_{T}\,p^{1}_{f,T}+o_{p}(1),\qquad\frac{1}{T}\sum_{t\in\mathcal{T}^{2}_{f,T}}\tilde{d}_{t,T}=\alpha_{T}\,p^{2}_{f,T}+o_{p}(1),$
where probability is over the random draw of $\sigma$ conditional on $(\mathcal{A}_{T},Z_{T})$. Plugging in gives
$M_{f,T}(\sigma)=\alpha_{T}\,p^{1}_{f,T}+(1-\alpha_{T})\,p^{2}_{f,T}+o_{p}(1).$ (A.6)
Since $p^{1}_{f,T}+p^{2}_{f,T}\leq 1$, we obtain the deterministic bound (for each $f$)
$\alpha_{T}\,p^{1}_{f,T}+(1-\alpha_{T})\,p^{2}_{f,T}\ \leq\ \max\{\alpha_{T},1-\alpha_{T}\}\,(p^{1}_{f,T}+p^{2}_{f,T})\ \leq\ \max\{\alpha_{T},1-\alpha_{T}\}.$
Combining with (A.6) and maximizing over $f$ yields
$m_{T}(\sigma):=\max_{f\in\mathcal{F}}M_{f,T}(\sigma)\ \leq\ \max\{\alpha_{T},1-\alpha_{T}\}+o_{p}(1).$ (A.7)
Finally, combine (A.5) and (A.7). Since the function $u\mapsto u^{2}+(1-u)^{2}$ is increasing on $[1/2,1]$ and $\max\{\alpha_{T},1-\alpha_{T}\}\in[1/2,1]$, we get
$T(D_{T,\sigma})\ \leq\ m_{T}(\sigma)^{2}+\big(1-m_{T}(\sigma)\big)^{2}\ \leq\ \alpha_{T}^{2}+(1-\alpha_{T})^{2}+o_{p}(1)\ =\ \tau_{0}(\alpha_{T})+o_{p}(1),$
with probability over $\sigma$ conditional on $(\mathcal{A}_{T},Z_{T})$.
Step 3 (Separation implies $p_{\mathrm{perm}}(D_{T})\to 0$). By hypothesis $\alpha_{T}\to_{p}\bar{\alpha}$ and $\tau(\pi)>\tau_{0}(\bar{\alpha})$, so there exists $\varepsilon>0$ such that $\tau(\pi)\geq\tau_{0}(\bar{\alpha})+3\varepsilon$.
By (A.2), $T(D_{T})\geq\tau(\pi)+o_{p}(1)$, hence
$\Pr\!\big(T(D_{T})\geq\tau_{0}(\bar{\alpha})+2\varepsilon\big)\ \longrightarrow\ 1.$
By Step 2 and $\alpha_{T}\to_{p}\bar{\alpha}$, conditional on $(\mathcal{A}_{T},Z_{T})$ the permutation draw satisfies
$\Pr_{\sigma}\!\big(T(D_{T,\sigma})\leq\tau_{0}(\bar{\alpha})+\varepsilon\big)\ \longrightarrow\ 1,$
where $\Pr_{\sigma}$ denotes probability over the uniform random permutation. Therefore the randomization tail probability
$p_{\mathrm{perm}}(D_{T})=\Pr_{\sigma}\!\big(T(D_{T,\sigma})\geq T(D_{T})\ \big|\ D_{T}\big)$
converges to 0 in probability, which proves $p_{\mathrm{perm}}(D_{T})\to_{p}0$. The power statement $\Pr(p_{\mathrm{perm}}(D_{T})\leq\eta)\to 1$ follows immediately. For the Monte Carlo version, $\widehat{p}_{\mathrm{perm}}(D_{T})\to_{p}p_{\mathrm{perm}}(D_{T})$ when $B_{T}\to\infty$, so $\widehat{p}_{\mathrm{perm}}(D_{T})\to_{p}0$ as well.
Appendix C Random Rule Model (RRM)
This appendix briefly sketches a “Random Rule Model” (RRM) formulation and shows how, in a simple case, it nests the random decision rule that underlies our permutation test. The goal is not to develop a full theory of RRMs - which would take us too far from the main focus of the paper - but rather to clarify the behavioral interpretation of the null hypothesis.
Definition
Fix a finite rule library $\mathcal{F}$ and a dominance preorder $\geq_{FSD}$ on $X$. In the main text, a deterministic rule profile is a sequence
$\mathbf{f}=(f_{t})_{t=1}^{T}\in\mathcal{F}^{T}$
assigning one rule to each observed menu $A_{t}=\{L_{t}^{1},L_{t}^{2}\}$. Given $\mathbf{f}$, the agent behaves as if she applies $f_{t}$ at menu $A_{t}$ and chooses the (perceived) dominant alternative.
A Random Rule Model replaces the deterministic profile by a random one. Formally, an RRM consists of:
  • a probability space $(\Omega,\mathcal{G},\mathbb{P})$, and
  • a stochastic process $H=(H_{t})_{t=1}^{T}$ defined on $(\Omega,\mathcal{G},\mathbb{P})$ with values in $\mathcal{F}^{T}$, where $H_{t}(\omega)\in\mathcal{F}$ is the rule applied at menu $A_{t}$ in state $\omega\in\Omega$.
such that, for each realization $\omega\in\Omega$, the agent behaves as if she were using the deterministic profile $H(\omega)=(H_{t}(\omega))_{t=1}^{T}$ menu by menu. The randomness in $H$ can be interpreted as unobserved heterogeneity in rule usage across time or across realizations of the decision maker. This construction parallels the Random Utility Model (McFaddenRichter1991; mcfadden2006) or the Random Attention Model of manzini_mariotti2014, with rules taking the place of utilities or attention filters.
In full generality, RRMs allow $H$ to have arbitrary dependence across $t$ and to depend on the menus $\{A_{t}\}_{t=1}^{T}$. In this paper, we do not attempt to characterize the class of choice distributions generated by such models. Instead, we focus on a simple special case that is directly linked to our random-choice null.
A simple RRM on attention rules
Consider the attention rules A1 and A2 introduced in the main text. Recall that at menu $A_{t}=\{L_{t}^{1},L_{t}^{2}\}$, rule A1 fully attends to (and selects) $L_{t}^{1}$, while A2 fully attends to (and selects) $L_{t}^{2}$. Let the heuristic process $H=(H_{t})_{t=1}^{T}$ take values in $\{\textsc{A1},\textsc{A2}\}$ and define the induced binary choice variables
$x_{t}=\begin{cases}1,&\text{if }H_{t}=\textsc{A1}\quad(\text{agent chooses }L_{t}^{1}),\\[1.99997pt] 0,&\text{if }H_{t}=\textsc{A2}\quad(\text{agent chooses }L_{t}^{2}).\end{cases}$
Under A1 the first lottery strictly dominates the second in the perceived menu, and under A2 the opposite holds. Hence, at each menu, the choice of which lottery is selected is equivalent to the choice of which attention rule is applied.
Let
$\alpha\ :=\ \frac{1}{T}\sum_{t=1}^{T}x_{t}$
be the observed fraction of $L_{t}^{1}$ choices. In the random-heuristic interpretation, $\alpha$ is simply the empirical share of menus at which the decision maker uses A1 rather than A2.
The next proposition shows that the random decision rule that underlies our permutation test in the main text can be viewed as a special case of an RRM on the library $\{\textsc{A1},\textsc{A2}\}$, and clarifies what the test conditions on.
Proposition A.2 (Random heuristic representation of the null).
Fix menus $\{A_{t}\}_{t=1}^{T}$. Suppose that under the null hypothesis $H_{0}$ the choices $(x_{1},\dots,x_{T})\in\{0,1\}^{T}$ are generated by a random decision rule that is exchangeable across $t$. Then:
  1. i
    There exists a Random Rule Model with rule library $\{\textsc{A1},\textsc{A2}\}$ and a sequence of random rules $(H_{t})_{t=1}^{T}$ such that
    $x_{t}=1\iff H_{t}=\textsc{A1},\qquad x_{t}=0\iff H_{t}=\textsc{A2}\quad\text{for all }t.$
    In particular, if $(x_{t})_{t=1}^{T}$ are i.i.d. $\mathrm{Bernoulli}(p)$, then $(H_{t})_{t=1}^{T}$ are i.i.d. with $\mathbb{P}(H_{t}=\textsc{A1})=p$, and $\alpha$ is the empirical share of attention to $L^{1}$.
  2. $i$
    Let $K:=\sum_{t=1}^{T}x_{t}$ be the realized number of $L_{t}^{1}$ choices (equivalently, of A1 uses). Conditional on $K$, the distribution of $(x_{1},\dots,x_{T})$ under $H_{0}$ is uniform over all binary vectors with exactly $K$ ones. Equivalently, the induced datasets
    $D_{\sigma}=\{(x_{\sigma(t)},A_{t})\}_{t=1}^{T},\qquad\sigma\in\Pi,$
    are equiprobable among all permutations of the $K$ ones and TKT-K zeros.
Proof.
For part (i), define $H_{t}:=\textsc{A1}$ whenever $x_{t}=1$ and $H_{t}:=\textsc{A2}$ whenever $x_{t}=0$. This yields a random sequence $(H_{t})_{t=1}^{T}$ taking values in $\{\textsc{A1},\textsc{A2}\}$ and satisfying $x_{t}=1\iff H_{t}=\textsc{A1}$ for every realization. If $(x_{t})$ are i.i.d. $\mathrm{Bernoulli}(p)$, then $(H_{t})$ are i.i.d. with $\mathbb{P}(H_{t}=\textsc{A1})=p$ by construction, and the statement about $\alpha$ follows.
For part (ii), exchangeability of $(x_{t})_{t=1}^{T}$ under $H_{0}$ means that for any permutation $\sigma$ of $\{1,\dots,T\}$ and any binary vector $x=(x_{t})_{t=1}^{T}$,
$\mathbb{P}\big((x_{1},\dots,x_{T})=x\big)=\mathbb{P}\big((x_{\sigma(1)},\dots,x_{\sigma(T)})=x\big).$
Conditional on $K=\sum_{t=1}^{T}x_{t}$, all binary vectors with exactly $K$ ones can be obtained from one another by permutations. Exchangeability therefore implies that they must receive the same conditional probability. This yields the uniformity statement and, in turn, the equiprobability of the permuted datasets $\{D_{\sigma}\}_{\sigma\in\Pi}$. ∎
Proposition A.2 shows that our permutation test evaluates the observed MRCI against a benchmark in which the agent follows a random rule on $\{\textsc{A1},\textsc{A2}\}$ with the same empirical frequencies $(\alpha,1-\alpha)$ of using each rule. The randomization distribution is therefore the exact conditional distribution of $T(D_{\sigma})$ under any exchangeable Random Rule Model of this form.
Appendix D Mixed Integer Quadratic Programming (MIQP)
This section develops a Mixed Integer Quadratic Programming (MIQP) approach to compute the exact value of the MRCI. We use this exact formulation to benchmark the accuracy and scalability of the heuristic optimization algorithm developed in Section 5.
D.1 MIQP Optimization
Precomputation.
Fix the rule library $\mathcal{F}$ and the dominance preorder $\geq_{\mathrm{FSD}}$. We define for every pair $(t,f)$ the strict admissibility indicator:
$a_{t,f}=\mathbf{1}\{L_{t}^{i}>_{\text{FSD}}L_{t}^{-i}\},$
where $L_{t}^{i}$ is the lottery chosen by rule $f$ at menu $t$, and $L_{t}^{-i}$ is the alternative.
Decision variable.
Let $y_{t,f}$ be a binary variable indicating whether rule $f$ is assigned to observation $t$:
$y_{t,f}\in\{0,1\}.$
Constraints.
We impose two sets of constraints to ensure a valid decomposition:
(R1) Partition Constraint. Each observation $t$ must be assigned to exactly one rule:
$\sum_{f\in\mathcal{F}}y_{t,f}=1\qquad\forall t.$ (R1)
(LD) Local Discriminability. A rule $f$ can only be assigned to observation $t$ if it strictly rationalizes the choice (i.e., strictly dominates the alternative according to the precomputed order):
$y_{t,f}\leq a_{t,f}\qquad\forall t,\ \forall f\in\mathcal{F}.$ (LD)
Proposition A.3.
The Mixed Integer Quadratic Program
[…]
computes the exact $\mathrm{MRCI}(D;\geq_{FSD},\mathcal{F})$.
D.2 Benchmarking: MIQP vs. Heuristic Approach
We compare the exact MIQP solution against the alternating maximization heuristic used in the main text (Section 5).
Methodology.
Since the MIQP problem is NP-hard, solving it for the full dataset ($T=750$ per subject) is computationally infeasible. To conduct a valid comparison, we generate subsampled datasets by randomly selecting $n\in\{1,2,3,4,5\}$ game trial from each of the 30 games for every subject. This yields a total of $3,430$ optimization tasks ranging in size from $T=30$ to $T=150$. For each task, we compute the MRCI using both the Gurobi MIQP solver and our heuristic algorithm with 100 random restarts, as in the main text. The Gurobi solver is free for academic use, and implemented via R.
Scalability.
Figure A.1 illustrates the computational necessity of the heuristic approach. As the sample size increases from 1 to 5 observations per game, the average runtime of the heuristic is $\approx 0.3$ seconds. In contrast, the runtime for the exact MIQP scales, reaching an average of 200 seconds for $n=5$. Extrapolating this to the full dataset ($n=25$) confirms that the exact method is intractable for large-scale analysis.
Figure A.1: Computational Scalability. The figure plots the average runtime (in seconds) required to compute the MRCI for a single subject as the number of sampled observations per game increases.
Accuracy.
Does the heuristic sacrifice accuracy for speed? We compute the absolute difference between the exact MRCI and the heuristic approximation: $\Delta=|\text{MRCI}_{\text{MIQP}}-\text{MRCI}_{\text{Heuristic}}|$. The results are summarized in Figure A.2.
The performance of the heuristic is near-exact. The mean absolute difference across all 2,058 tasks is $9.8\times 10^{-4}$ (approximately 0.001). Furthermore, 96.44% of all estimates fall within a strict 1% tolerance of the true global optimum. The maximum error observed in the entire benchmark was 0.085. These results provide strong evidence that the heuristic algorithm developed in Section 5 provides a high-quality approximation of the true MRCI while remaining computationally efficient for large datasets.
Figure A.2: Accuracy Benchmark. The figure displays the distribution of the absolute difference between the exact MRCI (computed via MIQP) and the heuristic approximation. The vertical dashed line indicates a 1% tolerance threshold.
Appendix E Additional Tables
Table A.1: OLS Regression of Rule Concentration Index (MRCI). * $p<0.1$, **$p<0.05$, *** $p<0.01$.
Dependent Variable: MRCI
(1) (2) (3)
Intercept 0.552∗∗∗ 0.550∗∗∗ 0.550∗∗∗
(0.003) (0.004) (0.004)
Female -0.013∗∗∗ -0.014∗∗∗ -0.014∗∗∗
(0.004) (0.004) (0.004)
Age -0.001 -0.001
(0.001) (0.001)
Age2 -0.000
Observations 686 686 686
R2 0.017 0.026 0.026
Figure A.3: Gender Heterogeneity in Rule Usage in subsample that passes the permutation test at the 1% level. The figure compares the Concentration Gain (Panel A) and Stability Score (Panel B) for men (dark blue) and women (orange).